### v3

# Further Comments on Yablo’s Construction ††thanks: File: WAB, [Sch25], arXiv 2504.10370

Abstract
We continue our analysis of Yablo’s coding of the liar paradox by infinite acyclic graphs, [Yab82].
The present notes are based on and continue the author’s previous results on the problem.
In particular, our approach is often more systematic than before.
Contents
1 Overview
These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions. All these papers are, of course, based on Yablo’s seminal [Yab82].
See also the author’s earlier versions on arXiv.
In Section 2 (page 2), we first present the background, its logical and graph theoretical side, and their interplay. We also discuss a useful third truth value, which expresses that neither nor is consistent.
In most cases, we will work with simple conjunctions (with and without negations).
In Section 3 (page 3), we discuss elementary contradiction (cells). Essentially, these are variations of Yablo’s triangles. The simplest contradictions are too simple, they have an “escape possibility”, i.e., combining them will leave one possibility without contradictions. Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework.
In Section 4 (page 4), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction, but give a finer analysis, discuss its properties, and again escape possibilities (in indirect composition).
We may start the construction with some simple but we will automatically end in a Yablo like construction, when satisfying the basic requirements (C1) and (C2) (see below, Section 1.1 (page 1.1)). This is discussed in the sections on “Saw Blades”, see Section 7.6 in [Sch22] and Section 3 in [Sch23b].
- (4)
1.1 Basic Terminology
Yablo’s construction
Yablo’s construction is an infinite directed graph with nodes and (negative) arrows for
Nodes stand for propositional variables, and the meaning of arrows is that variable occurs (negatively) in the formula attached to in most cases it expresses (We will identify for simplicity nodes with their attached variables.)
It is easy to see that no node may be true, nor false: Suppose is true, than all are false, so must be false, then contradiction, as all must be false. Suppose is false, then some must be true, again a contradiction (to the above).
Yablo Cell, Yablo Triangle
The basic contradiction structure has the form with the meaning (so and
We sometimes call the head, the knee, and the foot of the Yablo Cell (or Triangle).
Conditions (C1) and (C2)
We sometimes write when we want to emphasize that we assume is positive, similarly for the negative case.
Consider in the Yablo construction (but this applies to all
We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction. (More precisely, see Section 4.1 (page 4.1), (3).)
1.2 Comments On the Strategy
These comments will be clearer in hindsight, after having read the details. Still the best place to put them seems to be in the introduction.
The comments on “ is false ” show that there is no other way to express this in an acyclic graph of our type.
The comments on the construction in the limit solve a seeming contradiction by chosing the right level of abstraction, i.e. by arguing on the level of instances. We do not show that there cannot be any other way to do this - we do not know, we just see no other possibility.
1.2.1 Expressing “x is false” In an Acyclic Graph
is sometimes expressed by and by a contradiction.
How can we express that is wrong in a finite acyclic graph?
Consider the set of minimal elements below , and the formula corresponding to the graph in DNF of these elements. This might be
(the graph is trivial, just
or a tautology or contradiction.
a disjunction of conjunctions of elements in , or their negations. By acyclicity,
So the only way to express that a is wrong is by a contradiction.
1.2.2 Satisfying the Antagonistic Requirements (C1) and (C2) In the Limit
We want to express by an infinite graph that both and are inconsistent.
By (102.2), this must be expressed by suitable contradictions. Therefore, we treat such contradictions as basic building blocks, as Yablo did.
The problems are then
- (1)
How do we combine them to obtain a suitable graph expressing that both and are impossible?
The problem is that, if is contradictory, then is a tautology in classical logic, and thus in the theory expressed by a finite acyclic graph. So we alternate between and a tautology, and it is not clear what such constructions are in the limit, when we leave classical logic.
This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 4 (page 4).
We have to look at instances.
- (2.1)
- (2.2)
But is now a branching point. The branch continues to be satisfied for (C1), but the (new) branch opens a new problem for (C1), as is an OR for
- (2.4)
So we have new instances of problems for (C1), (C2), we solve the new instances, say for (C1), the old instances for both (C1) and (C2) stay solved.
Thus, in the limit, new problems (instances of problems, more precisely) are created, they are solved in one of the next steps, problems once solved stay solved, so there remain no unsolved problems.
The set of problems increases, but all are solved in the limit, and and are contradictory.
We have (essentially) an alternation of creating and treating problems for (C1) and (C2): We cannot treat e.g. a (C1) problem, bevor it was created by solving a (C2) problem - and conversely.
(Depending on organization, we have some liberty: e.g. solving (C1) creates a (C2) problem, but this needs TWO branches. So we might first treat these two branches ((C1) problems) before going back to the new (C2) problem created by the first new branch.)
2 Background: Graph, Logik
Most of the material in this section is either common knowledge, or was already covered in [Sch23b].
2.1 The Logical Side
On the logics side, we work with propositional formulas, which may, however, be infinite.
We will see that we need infinite formulas (of infinite depth) to have nodes to which we cannot attribute truth values. See Fact 2.4 (page 2.4).
We will work here with disjunctive normal forms, i.e. with formulas of the type where and the are propositional variables or negations thereof - most of the time pure conjunctions of negations of propositional variables.
Fact 2.1
Let where the are propositional variables or negations thereof.
Let - the set of choice functions in the sets where
Then
(Argue semantically with the sets of models and general distributivity and complements of sets.)
Contradictions will be between two formulas only, one a propositional variable, the other the negation of the former.
2.2 The Graph Side
We work with directed, acyclic graphs. They will usually have one root, often called In diagrams, the graphs may grow upward, downward, or sideways, we will say what is meant.
Definition 2.1
Nodes stand for propositional variables.
If a node is not terminal, it has also a propositional formula attached to it, sometimes written with the meaning abbreviated The successors of are the propositional variables occuring in Thus, if and are the only successors of in may be but not
Usually, the are (possibly infinite) conjunctions of propositional variables or (in most cases) their negations, which we write etc. We often indicate the negated variables in the graph with negated arrows, like etc. Thus, usually stands for
Example 2.1
in above definition cannot be replaced by etc, as this example shows (we argue semantically, with the central conflictual construction in Yablo’s paper, Yablo Cell, see Section 1.1 (page 1.1), (2)):
Let so so we have consistency, and it does not work for the construction.
Example 2.2
Consider the basic construction of a contradiction (used by Yablo and here, see Section 1.1 (page 1.1).
stands for so which is impossible.
If we negate then so is possible.
From the graph perspective, we have two paths in from to and
We add now to so thus so and and is not contradictory any more.
Definition 2.2
(This applies only to unique occurrences of a variable in the formula attached to another variable.)
We can attribute a value to a path expressing whether it changes a truth value from the beginning to the end. does not change the value of compared to that of does. We say or positive (negative) path.
Formally:
Let be paths as usual.
If then if then
Let be the concatenation of and Then iff and - otherwise.
If all arrows are negative, then iff the length of is even.
Definition 2.3
We call two paths with common start and end contradictory, and the pair a contradictory cell iff The structures considered here will be built with contradictory cells.
Remark 2.2
Note that the fact that and are contradictory or not is independent of how we start, whether for both or for both
We saw already in Example 2.2 (page 2.2) that it is not sufficient for a “real” contradiction to have two contradictory paths.
We need
(somewhere) an “AND”, so both have to be valid together, an “OR” is not sufficient,
we must not have a branching with an “OR” on the way as in an “escape” path, unless this leads again to a contradiction.
Notation 2.1
When we give nodes a truth value, we will use (and etc. if has the form to denote the case and etc. for the case
Fact 2.3
(Simplified).
Consider three paths, for simplicity with same origin, i.e.
No contradiction loops of length 3.
Suppose they meet at a common point, i.e. Then it is impossible that contradicts contradicts contradicts (at (“ contradicts ” means here that for some but one has value the other value -.)
(Trivial, we have only 2 truth values).
Suppose, first and meet, then (or and meet, but once they meet, they will continue the same way (e.g., if then for all ). Then it is again impossible that contradicts contradicts contradicts and continue to be the same but with different truth values until they meet so it is the same situation as above.)
Above properties generalize to any loops of odd length (with more paths).
This does not hold when the paths may branch again after meeting, as the next Example shows.
Example 2.3
(Example 7.4.2 in [Sch22].)
Let
then contradicts in the lower part, and in the upper part, contradicts and in the upper part, contradicts in the lower part.
Obviously, this may be generalized to paths.
Consider Yablo’s original construction:
Example 2.4
Let the nodes be and the arrows for expressed as a relation by and as a logical formula by
Thus For any we have a contradiction by and for any and for any for a suitable
2.3 Interplay of the Graph and the Logical Side
We can either think on the logical level with formulas, or graphically with conflicting paths, as in the following Fact.
We need infinite depth and width in our constructions:
Fact 2.4
The construction needs infinite depth,
the logic as used in Yablo’s construction is not compact,
the construction needs infinite width, i.e. the logic cannot be classical.
Proof: See Fact 1.4, 858 in [Sch23b].
2.4 A Third Truth Value
See also Section 7.1.2, Fact 7.1.1 in [Sch22] for a less systematic approach.
Remark 2.5
We consider a third truth value, in addition to and
means for a formula that neither nor are true (in the given set of formulas). Of course, this is impossible in a classical framework.
Let be the truth value of
Let be an abbreviation for and
Let
We have
analogously.
Likewise
We thus have:
and
Using the operations for truth values
and above results, we have
(or and thus the order
(and finally the inverse operations
Definition 2.4
We loosely use the expression “escape possibility”, “escape path” etc. to indicate a construction which avoids a desired property through a (usually infinite) sequence of decisions.
Examples:
A simple example is the trivial contradiction, so (See Definition 3.1 (page 3.1).) If we continue with so etc. we have an escape sequence etc., which never meets a contradiction.
Thus, is not a contradiction cell for our purposes, has a model defined by the escape sequence.
“Pipelines”: We may try to avoid infinite branching by one branch used to branch infinite often, but sequentially.
Say, instead of a disjunction of infinitely many arrows we construct one arrow and branch or or etc., but now the sequence …. avoids hitting any
In general, suppose we have a partial graph with origin and end point (among perhaps other end points), and, say results in choices or (due to internal choices by OR).
Remark 2.6
Suppose we have an escape problem, like thus
Can we append a new finite structure to which reduces the possibilities to just one value, say for So, whatever the input, or the outcome is
This, however, is not possible.
There are, whatever the inner structure, just four possibilities: it might be equivalent to The first two cases are trivial. Suppose Then Suppose Then
3 Elementary Cells
3.1 General Remarks
We try to describe here the basic constructions of contradictions.
As said in Fact 2.1 (page 2.1) (and re-written graph-theoretically), these consist of two branches with common origin, which meet again, and have different polarity. We call such constructions cells.
First, we want to exclude some trivialities. We describe only the part beginning at the branching point, not before, not after.
Second, we will define a simple hierarchy of such cells, and will allow a cell only the use of simpler cells as substructure. This prevents “cheating”. (See Remark 3.1 (page 3.1).) Thus, we look at those cells only which allow to construct a Yablo-like structure without the use of more complicate cells.
We first use (almost) only negative arrows, and nodes whose formulas are conjunctions. We will see how to generalize to more complicated paths.
Definition 3.1
We will argue that it suffices to consider these types of cells, see Remark 3.3 (page 3.3) and Section 3.2 (page 3.2). Furthermore, we can also neglect all but the (slightly generalized) Yablo Cells, as we will see in the present section.
It will become clear in a moment that above cells are fundamentally different, but for this we have to consider the second requirement (C2).
Remark 3.1
Suppose we add to the Yablo construction an additional node (or will have the same properties (i.e. truth value as but this addition will not contribute anything to the construction.
To prevent this, we require that it is possible to construct a Yablo structure using only the contradiction cell discussed (and perhaps simpler contradiction cells).
Remark 3.2
We recall again the requirement (C2), has to be contradictory. As (or and thus (or we add the requirement that and (or and are each contradictory.
Thus, Yablo cells and diamonds are different from the simple contradiction.
Yablo cells and diamonds are different from each other as long as we require in the latter case that the truth value of too is and not just False. We will see that making and leads to problems.
(Conversely, making in a Yablo cell not contradictory makes the construction, of course, equivalent to the simple contradiction.)
Remark 3.3
False, True
We present a systematic treatment of variants of the Yablo triangle (and the diamond).
In the Yablo construction, we attach to the Yablo triangle to and constructions, which make We also consider here the cases where we end by or and set e.g. etc., see (2.3.1) below.
We consider:
Do we have a contradiction for
Do we have contradictions for
Do we have escape possibilities?
Do we obtain
- (2.2.5)
We consider the cases in all possible combinations:
Case is equivalent to case
Thus, we look at e.g. case i.e. and etc., and consider whether the requirements in (2.2) above are satisfied.
We first look only at the minimal requirements
has to be contradictory, condition (2.2.1) above.
must not be classical condition (2.2.4) above.
The cases:
Consider so for any so these cases are classical.
Consider For is classical, so is, too. so Thus, is the only non-classical case among
Consider In cases so too. In case so In case so So is the only classical case among the
So the interesting cases left are
We now check for local contradiction in the case
This is simpler, we just have to consider the cases (e.g.) the are omitted, as we look for local contradictions, is anything, and, of course,
The case is known (Yablo construction, all cases except and
Case and we have no contradiction.
Case and we have a contradiction.
So are contradictory.
Next, we have to look at and see if every path from leads to a contradiction, condition (2.2.2) above.
We have to check it must not be
So and will not work, i.e.
Consider i.e. Then in and so this does not work. so this works. so this works.
Consider i.e. Then or and all cases work, except
In summary: in exactly the following cases:
We summarize the three conditions above:
(2.2.4): are ok.
(2.2.2): are not ok, so we learn nothing new beyond (2.2.4).
(2.2.1): are ok, so only is eliminated from those satisfying (2.2.4).
In Summary, only satisfy all three conditions, (2.2.1), (2.2.2), (2.2.4).
We look at these cases.
Here, and in (2.3.6), stands for which is of type likewise as we will append in the full structure at and a construction with value
In all cases
Thus,
Thus,
Thus,
identical to
Thus,
Thus,
We look at the Diamond.
Thus,
The modified Diamond is the same, only so and so we have a Yablo cell.
Summary of the (in this context) important constructions
3.2 Comments on Yablo Cells
Prerequisites:
We consider a triangle , , but leave open if the arrows are positive or negative. We require that is of type and that the triangle is contradictory for the case as well as for both paths und have to lead to a contradiction (conditions (C1) and (C2)).
3.2.1 Branching points
We consider the hierarchy:
intermediate point - branching point - branching point with contradiction
We have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).
If we have a contradiction on only at the situation is again equivalent to the trivial contradiction.
More general, if and contradict each other, and both are for x- valid paths (i.e. no contradiction at then we have an escape possibility.
The points
Suppose we branch at so we have, in addition to the triangle, some
Case 1: If then AND, thus if x- then OR:
Consider , , so, whatever the choice in we have what we need.
Without a contradiction at we have an escape possibility for (see above), e.g. continue the same way: , ,
Case 2: implies OR: (Yablo)
Thus, we also need for a contradiction, this leads to infinite branching at
A new path as in Yablo’s construction leads to infinite branching and non-classical logic.
We may branch on the already existing oder , and continue to build a (finite) contradiction to
This however, is possible only a finite number of times, e.g. then etc., like “Pipeline”, Example 7.37, and Section 7.6.3 in [Sch22] or Example 1.5 and Section 3.4 in [Sch23b]. See also Definition 2.4 (page 2.4) above. This constructs an infinite sequence of choices which offers an escape possibility.
Thus we need infinite branching (and non-classical logic).
We see that we may consider only the new paths they form an infinite subsequence of the construction. By full transitivity of the final construction, any infinite subsequence of the construction has the same properties as the full construction.
3.2.2 Paths
- (1)
We branch on , and have , ,
Case 1: OR:
We make a copy of the structure for
Case 2: AND:
So x- OR:
We make a copy of the structure for
Consider e.g. , ,
As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (” pipeline “), we may assume for simplicity that we always start contradictions at
Diagram 3.1
3.2.3 All Sides Of the Triangle Have to Be Negative
Consider , . Suppose all are of the type
To obtain a contradiction for the triangle needs one or three negative sides.
Assume it has only one negative side.
Let , be a contradiction for (condition (C1)) and , be a contradiction for (for condition (C2)).
Suppose is positive, so we have or and, analogously, or (they have to be of the same type).
We leave aside the condition between and it is only the branching at we are interested in.
Thus, we have or
In the first case so and both and are possibilities, and, as in the simplest contradiction for we have an escape situation (see Definition 2.4 (page 2.4), (1)) (” watered down “by the additional possibilities and
In the second case so and the situation is analogous to the first case.
Suppose is negative, so we have and analogously
Now, we have to contradict for the case but for this we need a third negative arrow so this is impossible, and leads back to the usual Yablo cell.
There is a way out. Instead of considering we continue by This is equivalent to the Yablo cell by
But this is “cheating”.
3.2.4 The Problem with Diamonds
Remark 3.4
We have a conflict between the diamonds starting at and and the diamond starting at
If then y- and As the choices at and are independent, any branch -z, combined with any branch -z, must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
However, if is positive, the diamond -z, has to be contradictory, so not both branches may be negative.
A solution is to “synchronise” the choices at and which can be done e.g. by the formula
and This is a different type of formula, moreoever corresponding arrows from are missing. Even if we do not consider the paths for the diamond starting at we have and so we can make and false, and chose and which results in an escape possibility (we have to consider new diamonds starting at etc.)
This formula has an additional flaw, we want to speak about paths we chose, and not just their end points. This can be done with introducing additional points on the arrows, e.g. the arrow is replaced by now we can speak about the arrow we have given it a name.
We may also put such information in a background theory, which need no be negated.
But all this leads too far from the basically simple formalism of Yablo’s construction, so we will not discuss this any further.
(Yablo’s construction does not need synchronisation, this is done automatically by the universal quantifier.)
The simplified version with recursion only on the left is nothing but the Yablo triangle:
Thus, we conclude that the original version with 4 points and full recursion is beyond our scope, and the simplified version (see above) is nothing but the Yablo triangle.
Diagram 3.2
Nested Diamonds, Details
Diagram 3.3
Diamond Essentials, Recursion on the Left, Simple Contradiction on the Right
4 Combining Cells
4.1 Overview
We follow Yablo’s construction.
(C1): We have to make contradictory.
We first try to make “directly” contradictory. That is, we will not first go to and make contradictory, nor will we make a disjunction at and then make each disjunct contradictory.
(This is not an important simplification, as more complicated constructions will finally use our simple one, too. See e.g. Diagram 4.3 (page 4.3) and its discussion.)
As Yablo did, we set and
The first step is done, is contradictory. Probabilistically, of the possibilities via are contradictory.
(C2) We have to make contradictory.
Now, must be contradictory, too. As we do not know which one is positive, we must make both and contradictory - just as a disjunction is false if every disjunct is false.
Consider, e.g., Again, we will make directly contradictory (as above for so is a conjunction, say (we might re-use here, but this will not be important) and add e.g. and to for the contradiction, so we have
But now, we have destroyed the contradiction at as for is a disjunction, and have to make ALL possibilities for contradictory for so far only the possibility is contradictory.
Probabilistically, of the possibilities at are open.
To create a contradiction for the possibility we introduce new contradictions by adding to so
Etc.
See below and Section 3.3.3. in [Sch23b].
Note that for (C1), i.e. for and for (C2), i.e. for we have to make for any disjunction all possibilities contradictory, and for any conjunction it suffices to make one conjunct contradictory, as and
Of course, following a path in the graph, conjunctions for will become disjunctions for one of the confusing aspects of reasoning.
(C1) and (C2) are antagonistic requirements. Satisfying (C1) creates new problems for (C2), and vice versa. In the limit, all requirements are satisfied, since all problems will solved in the next step, at the price of creating new problems, which will be satisfied in the next step.
Note that we don’t approximate, as long we are finite, we always have only in the infinite case we have
The right level of abstraction:
The author tried first to work towards a representation result using single arrows. However, the following result suggests that one should work with (perhaps composite) paths instead of arrows: a single negative arrow may be replaced by the following diagram:
So
Thus, we may replace negative arrows recursively by arbitrarily deep constructions using negative arrows again, complicating the same construction arbitrarily.
The same is true for all other logical operators.
(See Section 4 in Sch23b.)
4.2 In More Detail
We now consider the inductive construction of the Yablo structure, see Diagram 4.1 (page 4.1) and Diagram 4.2 (page 4.2)
In (2) and (5), we have unrelated points and we may consider the model sets to be orthogonal. - We only indicate model sets briefly, without going into details (U will stand for the universe).
abbreviated
has to lead to a contradiction for
preparation,
was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
finish,
(branch has to be contradicted for add
has to lead to a contradiction for
preparation,
finish,
(branch has to be contradicted for add
(branch has to be contradicted for add
Construction 4.1
Summary:
Thus, there are arrows for all and the construction is transitive for the
Every is head of a Yablo cell with knee
Thus, every arrow from any to any goes to the head of a Yablo cell, and not only the arrows from
This property is “accidental”, and due to the fact that for any arrow there is also an arrow and property (2) holds for by prerequisite.
The construction has infinite depth and branching.
Diagram 4.1
Diagram Inductive Construction 1
Diagram 4.2
Diagram Inductive Construction 2
4.2.1 Composition
The prerequisites for the constructions in the diagrams Diagram 4.1 (page 4.1) etc. are negative Paths, und formulas.
These prerequisites hold also in our generalization to paths (modulo some simplifications for origins at etc.), so these constructions are also valid in the generalization to paths.
Remark 4.1
The condition “Existence of negative paths” is not trivial. Suppose that and are both negative, and is the only path from to then the condition is obviously false. The author does not know how to characterize graphs which satisfy the condition. Some kind of “richness” for the graph will probably have to hold. (This is covered in Yablo’s construction by transitivity.)
Example 4.1
There is a fundamental difference between the contradictions (for and
Consider the case The path is blocked at So we cannot go to by two contradicting possibilities. The path is, however, not blocked at (or elsewhere), so we have two contradicting possibilities to go from to This basically behaves like and we have an escape possibility. So, if we try to append the same construction at we have an escape possibility, always making the origin of the construction negative. This is not true for We can append the same construction at as will always be at the opposite polarity of due to , the alternative path ends at so there is no escape. So a negative start at the original diagram leads to a positive start in the appended diagram.
Diagram 4.3
Diagram Postponed Contradictions
4.3 Beyond Conjunctions
We work with disjunctive normal forms.
Consider
As has to be contradictory, both components and have to be contradictory.
As has to be contradictory, all components, etc., have to be contradictory. For this, it is sufficient and necessary that in each component, one element is contradictory. This is equivalent to the condition that (both and or (both and are contradictory.
Thus
For both
and
have to be contradictory, this is done as usual by
and
For we have to attach a contradiction at either
(-1) and
or
(-2) and
We thus have to examine the combinations
(-1)
(-2)
(-1)
(-2)
The combinations (-1)) and (-2)) are as usual, the other ones are new.
It seems that there might be a problem.
In the cases and seems to be independent from and both solvable by classical logic, a contradiction. But all cases are parts of and and (-1)) and (-2)) need simultaneous treatment of the and the - part.
So the confusion is only transient.
(In the case we have no independent parts, so the “problem” just discussed will not occur.)
5 Conjecture
Consider structures of the type (like ’s), with one small addition: we also allow formulas of the type either as (stand alone), or as one of several elements of a conjunction where is of the type
(This is more general than adding the constant TRUE, as it permits chaining of the new formulas, but we also add TRUE for terminal
In addition, we assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it. We also assume we use only one type of contradiction cell, and work not first with unsuitable cells, postponing the use of efficient cells. This covers also the “interdiction” of escape paths. See here e.g. the beginning of Section 3.1 (page 3.1), and Remark 3.1 (page 3.1), Definition 2.4 (page 2.4), and Example 4.1 (page 4.1), with Diagram 4.3 (page 4.3). Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
5.1 Conjecture: A Suitable Injection Does the Job of Characterisation
Consider Yablo’s original structure YS, and some directed acyclic graph YS’.
Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory iff
There is an injection from YS to YS’, more precisely from the nodes of YS to the nodes of YS’, s.t., if there is a path from to in YS, then there is a path from to
For THIS injection, there is a valuation of the arrows in (the other arrows are not interesting) s.t., if form a contradiction in YS, then likewise for (Contradictions are preserved under
5.2 Idea of Proof (Outline)
Fact 5.2
Yablo’s structure is the simplest structure showing that both and are contradictory.
Proof
Earlier sections of this paper. More precisely:
- (1.1)
- (1.2)
- (1.3)
Remark 5.3
Consider
We may chain such formulas together, like etc., which is impossible for the constant TRUE.
is preserved, if then too. Likewise, if and then
Consider Yablo adding
EG: so , this works well classically, but also for as
We may also continue beyond hinaus: or then too.
If we have mixed chains of and the last element will decide the result (trivial).
Fact 5.4
Adding new formulas inside a Yablo Triangle does not improve anything, it only destroys. It leads either to no contradiction, or to a contradiction of the simplest type, and thus to an escape situation, see Definition 2.4 (page 2.4).
There are no other new contradictory pairs of paths.
Proof
Consider a YC
Insert in ab, or and So, if then if then
Consider now insertions of xy, or xz, yz (analogously) and arbitrary combinations thereof.
xz: If then of then so we have a trivial contradiction and an escape possibility.
xy: xy as so If then if then we have again a trivial contradiction and an escape possibility.
yz: so If then if then we have again a trivial contradiction and an escape possibility.
Combinations:
For combinations of xy and yz, see (2.5) above.
xy and xz: the result is TRUE (both positive) or trivial contradiction with ESC.
xz and yz: This is analogous to xy and xz.
Let be such paths originating in containing some xy as in Fact 5.4 (page 5.4). Then the same considerations as above apply, so we have either no contradiction, or the trivial one with an escape possibility.
Similar results hold when using the constant TRUE: If has -TRUE in one branch, then will have TRUE in this branch, violating (C2). If has TRUE in one branch, then this branch is irrelevant, and we have no contradiction. So TRUE must not occur.
Extend the interpretation beyond
Let If has no successor, then set
If has successors etc., then set
If then has successors. Take the conjunction of the following: if if then
We have to show that still does what it should.
During construction, will be and after completion, And any will be TRUE or By the combination with AND, will not change.

### v4

@@
 Contents
-1 Overview
-These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions. All these papers are, of course, based on Yablo’s seminal [Yab82].
+1 Introduction
+Please note that arXiv pdf files sometimes have problems with some of the diagrams. In that case, please consult the html version.
+1.1 Outline
+These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions.
 See also the author’s earlier versions on arXiv.
-In Section 2 (page 2), we first present the background, its logical and graph theoretical side, and their interplay.
+In particular, the reader is referred there for a discussion of more complicated formulas attached to the nodes.
+We liberally use material from these sources.
+All these papers are, of course, based on Yablo’s seminal [Yab82].
+In Section 1.2 (page 1.2), we first present the background, its logical and graph theoretical side, and their interplay.
 We also discuss a useful third truth value, which expresses that neither nor is consistent.
 In most cases, we will work with simple conjunctions (with and without negations).
-In Section 3 (page 3), we discuss elementary contradiction (cells).
+In Section 2 (page 2), we discuss elementary contradictions (cells).
 Essentially, these are variations of Yablo’s triangles.
@@
 Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework.
-In Section 4 (page 4), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox.
-Basically, we follow Yablo’s construction, but give a finer analysis, discuss its properties, and again escape possibilities (in indirect composition).
-We may start the construction with some simple but we will automatically end in a Yablo like construction, when satisfying the basic requirements (C1) and (C2) (see below, Section 1.1 (page 1.1)).
-This is discussed in the sections on “Saw Blades”, see Section 7.6 in [Sch22] and Section 3 in [Sch23b].
+In Section 3 (page 3), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction.
+We then generalize slightly, see Remark 3.1 (page 3.1), but show that this general construction contains Yablo’s construction, so it is minimal.
+(See also the discussion of “Saw Blades” in Section 7.6 of [Sch22] and Section 3 in [Sch23b].)
+In both cases, we show that - within the chosen variant - the construction is necassary.
 - (4)
-1.1 Basic Terminology
+There is an important property for the graphs, which we need, probably a sort of “richness”, see Remark 1.5 (page 1.5).
+This property holds trivially in Yablo’s construction by transitivity.
+In our approach, it will hold by the construction of a suitable injection, see Section 4 (page 4), Condition 4.1 (page 4.1) (2).
+1.2 Basic Definitions and Facts
+1.2.1 Terminology
 Yablo’s construction
@@
 We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction.
-(More precisely, see Section 4.1 (page 4.1), (3).)
-1.2 Comments On the Strategy
-These comments will be clearer in hindsight, after having read the details. Still the best place to put them seems to be in the introduction.
-The comments on “ is false ” show that there is no other way to express this in an acyclic graph of our type.
-The comments on the construction in the limit solve a seeming contradiction by chosing the right level of abstraction, i.e.
-by arguing on the level of instances.
-We do not show that there cannot be any other way to do this - we do not know, we just see no other possibility.
-1.2.1 Expressing “x is false” In an Acyclic Graph
-is sometimes expressed by and by a contradiction.
-How can we express that is wrong in a finite acyclic graph?
-Consider the set of minimal elements below , and the formula corresponding to the graph in DNF of these elements. This might be
-(the graph is trivial, just
-or a tautology or contradiction.
-a disjunction of conjunctions of elements in , or their negations. By acyclicity,
-So the only way to express that a is wrong is by a contradiction.
-1.2.2 Satisfying the Antagonistic Requirements (C1) and (C2) In the Limit
-We want to express by an infinite graph that both and are inconsistent.
-By (102.2), this must be expressed by suitable contradictions. Therefore, we treat such contradictions as basic building blocks, as Yablo did.
-The problems are then
-- (1)
-How do we combine them to obtain a suitable graph expressing that both and are impossible?
-The problem is that, if is contradictory, then is a tautology in classical logic, and thus in the theory expressed by a finite acyclic graph.
-So we alternate between and a tautology, and it is not clear what such constructions are in the limit, when we leave classical logic.
-This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 4 (page 4).
-We have to look at instances.
-- (2.1)
-- (2.2)
-But is now a branching point. The branch continues to be satisfied for (C1), but the (new) branch opens a new problem for (C1), as is an OR for
-- (2.4)
-So we have new instances of problems for (C1), (C2), we solve the new instances, say for (C1), the old instances for both (C1) and (C2) stay solved.
-Thus, in the limit, new problems (instances of problems, more precisely) are created, they are solved in one of the next steps, problems once solved stay solved, so there remain no unsolved problems.
-The set of problems increases, but all are solved in the limit, and and are contradictory.
-We have (essentially) an alternation of creating and treating problems for (C1) and (C2): We cannot treat e.g. a (C1) problem, bevor it was created by solving a (C2) problem - and conversely.
-(Depending on organization, we have some liberty: e.g.
-solving (C1) creates a (C2) problem, but this needs TWO branches.
-So we might first treat these two branches ((C1) problems) before going back to the new (C2) problem created by the first new branch.)
-2 Background: Graph, Logik
-Most of the material in this section is either common knowledge, or was already covered in [Sch23b].
-2.1 The Logical Side
+(More precisely, see Section 3.1 (page 3.1), (3).)
+1.2.2 The Logical Side
 On the logics side, we work with propositional formulas, which may, however, be infinite.
-We will see that we need infinite formulas (of infinite depth) to have nodes to which we cannot attribute truth values. See Fact 2.4 (page 2.4).
+We will see that we need infinite formulas (of infinite depth) to have nodes to which we cannot attribute truth values. See Fact 1.6 (page 1.6).
+Notation 1.1
+When we give nodes a truth value, we will use (and etc. if has the form to denote the case and etc. for the case
 We will work here with disjunctive normal forms, i.e.
 with formulas of the type where and the are propositional variables or negations thereof - most of the time pure conjunctions of negations of propositional variables.
-Fact 2.1
+Fact 1.1
 Let where the are propositional variables or negations thereof.
@@
 Contradictions will be between two formulas only, one a propositional variable, the other the negation of the former.
-2.2 The Graph Side
-We work with directed, acyclic graphs. They will usually have one root, often called In diagrams, the graphs may grow upward, downward, or sideways, we will say what is meant.
-Definition 2.1
-Nodes stand for propositional variables.
-If a node is not terminal, it has also a propositional formula attached to it, sometimes written with the meaning abbreviated The successors of are the propositional variables occuring in Thus, if and are the only successors of in may be but not
-Usually, the are (possibly infinite) conjunctions of propositional variables or (in most cases) their negations, which we write etc.
-We often indicate the negated variables in the graph with negated arrows, like etc.
-Thus, usually stands for
+Fact 1.2
+Let and we want to show that is contradictory, then it suffices to show that a subset of the is inconsistent - this corresponds to the fact that (X any truth value.)
+Let and we want to show that is contradictory, then we have to show that ALL are contradictory. Again, this corresponds to
+The following example illustrates the propagation of required contradictions.
 Example 1.1
-in above definition cannot be replaced by etc, as this example shows (we argue semantically, with the central conflictual construction in Yablo’s paper, Yablo Cell, see Section 1.1 (page 1.1), (2)):
-Let so so we have consistency, and it does not work for the construction.
-Example 2.2
-Consider the basic construction of a contradiction (used by Yablo and here, see Section 1.1 (page 1.1).
-stands for so which is impossible.
-If we negate then so is possible.
-From the graph perspective, we have two paths in from to and
-We add now to so thus so and and is not contradictory any more.
-Definition 2.2
-(This applies only to unique occurrences of a variable in the formula attached to another variable.)
-We can attribute a value to a path expressing whether it changes a truth value from the beginning to the end. does not change the value of compared to that of does. We say or positive (negative) path.
-Formally:
-Let be paths as usual.
-If then if then
-Let be the concatenation of and Then iff and - otherwise.
-If all arrows are negative, then iff the length of is even.
-Definition 2.3
-We call two paths with common start and end contradictory, and the pair a contradictory cell iff The structures considered here will be built with contradictory cells.
-Remark 2.2
-Note that the fact that and are contradictory or not is independent of how we start, whether for both or for both
-We saw already in Example 2.2 (page 2.2) that it is not sufficient for a “real” contradiction to have two contradictory paths.
-We need
-(somewhere) an “AND”, so both have to be valid together, an “OR” is not sufficient,
-we must not have a branching with an “OR” on the way as in an “escape” path, unless this leads again to a contradiction.
-Notation 2.1
-When we give nodes a truth value, we will use (and etc. if has the form to denote the case and etc. for the case
-Fact 2.3
-(Simplified).
-Consider three paths, for simplicity with same origin, i.e.
-No contradiction loops of length 3.
-Suppose they meet at a common point, i.e. Then it is impossible that contradicts contradicts contradicts (at (“ contradicts ” means here that for some but one has value the other value -.)
-(Trivial, we have only 2 truth values).
-Suppose, first and meet, then (or and meet, but once they meet, they will continue the same way (e.g., if then for all ).
-Then it is again impossible that contradicts contradicts contradicts and continue to be the same but with different truth values until they meet so it is the same situation as above.)
-Above properties generalize to any loops of odd length (with more paths).
-This does not hold when the paths may branch again after meeting, as the next Example shows.
-Example 2.3
-(Example 7.4.2 in [Sch22].)
-Let
-then contradicts in the lower part, and in the upper part, contradicts and in the upper part, contradicts in the lower part.
-Obviously, this may be generalized to paths.
-Consider Yablo’s original construction:
-Example 2.4
-Let the nodes be and the arrows for expressed as a relation by and as a logical formula by
-Thus For any we have a contradiction by and for any and for any for a suitable
-2.3 Interplay of the Graph and the Logical Side
-We can either think on the logical level with formulas, or graphically with conflicting paths, as in the following Fact.
-We need infinite depth and width in our constructions:
-Fact 2.4
-The construction needs infinite depth,
-the logic as used in Yablo’s construction is not compact,
-the construction needs infinite width, i.e. the logic cannot be classical.
-Proof: See Fact 1.4, 858 in [Sch23b].
-2.4 A Third Truth Value
-See also Section 7.1.2, Fact 7.1.1 in [Sch22] for a less systematic approach.
-Remark 2.5
-We consider a third truth value, in addition to and
-means for a formula that neither nor are true (in the given set of formulas). Of course, this is impossible in a classical framework.
-Let be the truth value of
-Let be an abbreviation for and
-Let
-We have
-analogously.
-Likewise
-We thus have:
-and
-Using the operations for truth values
-and above results, we have
-(or and thus the order
-(and finally the inverse operations
-Definition 2.4
-We loosely use the expression “escape possibility”, “escape path” etc. to indicate a construction which avoids a desired property through a (usually infinite) sequence of decisions.
-Examples:
-A simple example is the trivial contradiction, so (See Definition 3.1 (page 3.1).) If we continue with so etc. we have an escape sequence etc., which never meets a contradiction.
-Thus, is not a contradiction cell for our purposes, has a model defined by the escape sequence.
-“Pipelines”: We may try to avoid infinite branching by one branch used to branch infinite often, but sequentially.
-Say, instead of a disjunction of infinitely many arrows we construct one arrow and branch or or etc., but now the sequence …. avoids hitting any
-In general, suppose we have a partial graph with origin and end point (among perhaps other end points), and, say results in choices or (due to internal choices by OR).
-Remark 2.6
-Suppose we have an escape problem, like thus
-Can we append a new finite structure to which reduces the possibilities to just one value, say for So, whatever the input, or the outcome is
-This, however, is not possible.
-There are, whatever the inner structure, just four possibilities: it might be equivalent to The first two cases are trivial. Suppose Then Suppose Then
-3 Elementary Cells
-3.1 General Remarks
-We try to describe here the basic constructions of contradictions.
-As said in Fact 2.1 (page 2.1) (and re-written graph-theoretically), these consist of two branches with common origin, which meet again, and have different polarity. We call such constructions cells.
-First, we want to exclude some trivialities. We describe only the part beginning at the branching point, not before, not after.
-Second, we will define a simple hierarchy of such cells, and will allow a cell only the use of simpler cells as substructure.
-This prevents “cheating”.
-(See Remark 3.1 (page 3.1).) Thus, we look at those cells only which allow to construct a Yablo-like structure without the use of more complicate cells.
-We first use (almost) only negative arrows, and nodes whose formulas are conjunctions. We will see how to generalize to more complicated paths.
-Definition 3.1
-We will argue that it suffices to consider these types of cells, see Remark 3.3 (page 3.3) and Section 3.2 (page 3.2).
-Furthermore, we can also neglect all but the (slightly generalized) Yablo Cells, as we will see in the present section.
-It will become clear in a moment that above cells are fundamentally different, but for this we have to consider the second requirement (C2).
-Remark 3.1
-Suppose we add to the Yablo construction an additional node (or will have the same properties (i.e. truth value as but this addition will not contribute anything to the construction.
-To prevent this, we require that it is possible to construct a Yablo structure using only the contradiction cell discussed (and perhaps simpler contradiction cells).
-Remark 3.2
-We recall again the requirement (C2), has to be contradictory. As (or and thus (or we add the requirement that and (or and are each contradictory.
-Thus, Yablo cells and diamonds are different from the simple contradiction.
-Yablo cells and diamonds are different from each other as long as we require in the latter case that the truth value of too is and not just False. We will see that making and leads to problems.
-(Conversely, making in a Yablo cell not contradictory makes the construction, of course, equivalent to the simple contradiction.)
-Remark 3.3
-False, True
-We present a systematic treatment of variants of the Yablo triangle (and the diamond).
-In the Yablo construction, we attach to the Yablo triangle to and constructions, which make We also consider here the cases where we end by or and set e.g. etc., see (2.3.1) below.
-We consider:
-Do we have a contradiction for
-Do we have contradictions for
-Do we have escape possibilities?
-Do we obtain
-- (2.2.5)
-We consider the cases in all possible combinations:
-Case is equivalent to case
-Thus, we look at e.g. case i.e. and etc., and consider whether the requirements in (2.2) above are satisfied.
-We first look only at the minimal requirements
-has to be contradictory, condition (2.2.1) above.
-must not be classical condition (2.2.4) above.
-The cases:
-Consider so for any so these cases are classical.
-Consider For is classical, so is, too. so Thus, is the only non-classical case among
-Consider In cases so too. In case so In case so So is the only classical case among the
-So the interesting cases left are
-We now check for local contradiction in the case
-This is simpler, we just have to consider the cases (e.g.) the are omitted, as we look for local contradictions, is anything, and, of course,
-The case is known (Yablo construction, all cases except and
-Case and we have no contradiction.
-Case and we have a contradiction.
-So are contradictory.
-Next, we have to look at and see if every path from leads to a contradiction, condition (2.2.2) above.
-We have to check it must not be
-So and will not work, i.e.
-Consider i.e. Then in and so this does not work. so this works. so this works.
-Consider i.e. Then or and all cases work, except
-In summary: in exactly the following cases:
-We summarize the three conditions above:
-(2.2.4): are ok.
-(2.2.2): are not ok, so we learn nothing new beyond (2.2.4).
-(2.2.1): are ok, so only is eliminated from those satisfying (2.2.4).
-In Summary, only satisfy all three conditions, (2.2.1), (2.2.2), (2.2.4).
-We look at these cases.
-Here, and in (2.3.6), stands for which is of type likewise as we will append in the full structure at and a construction with value
-In all cases
-Thus,
-Thus,
-Thus,
-identical to
-Thus,
-Thus,
-We look at the Diamond.
-Thus,
-The modified Diamond is the same, only so and so we have a Yablo cell.
-Summary of the (in this context) important constructions
-3.2 Comments on Yablo Cells
-Prerequisites:
-We consider a triangle , , but leave open if the arrows are positive or negative.
-We require that is of type and that the triangle is contradictory for the case as well as for both paths und have to lead to a contradiction (conditions (C1) and (C2)).
-3.2.1 Branching points
-We consider the hierarchy:
-intermediate point - branching point - branching point with contradiction
-We have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).
-If we have a contradiction on only at the situation is again equivalent to the trivial contradiction.
-More general, if and contradict each other, and both are for x- valid paths (i.e. no contradiction at then we have an escape possibility.
-The points
-Suppose we branch at so we have, in addition to the triangle, some
-Case 1: If then AND, thus if x- then OR:
-Consider , , so, whatever the choice in we have what we need.
-Without a contradiction at we have an escape possibility for (see above), e.g. continue the same way: , ,
-Case 2: implies OR: (Yablo)
-Thus, we also need for a contradiction, this leads to infinite branching at
-A new path as in Yablo’s construction leads to infinite branching and non-classical logic.
-We may branch on the already existing oder , and continue to build a (finite) contradiction to
-This however, is possible only a finite number of times, e.g.
-then etc., like “Pipeline”, Example 7.37, and Section 7.6.3 in [Sch22] or Example 1.5 and Section 3.4 in [Sch23b].
-See also Definition 2.4 (page 2.4) above.
-This constructs an infinite sequence of choices which offers an escape possibility.
-Thus we need infinite branching (and non-classical logic).
-We see that we may consider only the new paths they form an infinite subsequence of the construction.
-By full transitivity of the final construction, any infinite subsequence of the construction has the same properties as the full construction.
-3.2.2 Paths
-- (1)
-We branch on , and have , ,
-Case 1: OR:
-We make a copy of the structure for
-Case 2: AND:
-So x- OR:
-We make a copy of the structure for
-Consider e.g. , ,
-As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (” pipeline “), we may assume for simplicity that we always start contradictions at
-Diagram 3.1
-3.2.3 All Sides Of the Triangle Have to Be Negative
-Consider , . Suppose all are of the type
-To obtain a contradiction for the triangle needs one or three negative sides.
-Assume it has only one negative side.
-Let , be a contradiction for (condition (C1)) and , be a contradiction for (for condition (C2)).
-Suppose is positive, so we have or and, analogously, or (they have to be of the same type).
-We leave aside the condition between and it is only the branching at we are interested in.
-Thus, we have or
-In the first case so and both and are possibilities, and, as in the simplest contradiction for we have an escape situation (see Definition 2.4 (page 2.4), (1)) (” watered down “by the additional possibilities and
-In the second case so and the situation is analogous to the first case.
-Suppose is negative, so we have and analogously
-Now, we have to contradict for the case but for this we need a third negative arrow so this is impossible, and leads back to the usual Yablo cell.
-There is a way out. Instead of considering we continue by This is equivalent to the Yablo cell by
-But this is “cheating”.
-3.2.4 The Problem with Diamonds
-Remark 3.4
-We have a conflict between the diamonds starting at and and the diamond starting at
-If then y- and As the choices at and are independent, any branch -z, combined with any branch -z, must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
-However, if is positive, the diamond -z, has to be contradictory, so not both branches may be negative.
-A solution is to “synchronise” the choices at and which can be done e.g. by the formula
-and This is a different type of formula, moreoever corresponding arrows from are missing.
-Even if we do not consider the paths for the diamond starting at we have and so we can make and false, and chose and which results in an escape possibility (we have to consider new diamonds starting at etc.)
-This formula has an additional flaw, we want to speak about paths we chose, and not just their end points.
-This can be done with introducing additional points on the arrows, e.g.
-the arrow is replaced by now we can speak about the arrow we have given it a name.
-We may also put such information in a background theory, which need no be negated.
-But all this leads too far from the basically simple formalism of Yablo’s construction, so we will not discuss this any further.
-(Yablo’s construction does not need synchronisation, this is done automatically by the universal quantifier.)
-The simplified version with recursion only on the left is nothing but the Yablo triangle:
-Thus, we conclude that the original version with 4 points and full recursion is beyond our scope, and the simplified version (see above) is nothing but the Yablo triangle.
-Diagram 3.2
-Nested Diamonds, Details
-Diagram 3.3
-Diamond Essentials, Recursion on the Left, Simple Contradiction on the Right
-4 Combining Cells
-4.1 Overview
-We follow Yablo’s construction.
-(C1): We have to make contradictory.
-We first try to make “directly” contradictory. That is, we will not first go to and make contradictory, nor will we make a disjunction at and then make each disjunct contradictory.
-(This is not an important simplification, as more complicated constructions will finally use our simple one, too. See e.g. Diagram 4.3 (page 4.3) and its discussion.)
-As Yablo did, we set and
-The first step is done, is contradictory. Probabilistically, of the possibilities via are contradictory.
-(C2) We have to make contradictory.
-Now, must be contradictory, too. As we do not know which one is positive, we must make both and contradictory - just as a disjunction is false if every disjunct is false.
-Consider, e.g., Again, we will make directly contradictory (as above for so is a conjunction, say (we might re-use here, but this will not be important) and add e.g.
-and to for the contradiction, so we have
-But now, we have destroyed the contradiction at as for is a disjunction, and have to make ALL possibilities for contradictory for so far only the possibility is contradictory.
-Probabilistically, of the possibilities at are open.
-To create a contradiction for the possibility we introduce new contradictions by adding to so
-Etc.
-See below and Section 3.3.3. in [Sch23b].
-Note that for (C1), i.e. for and for (C2), i.e. for we have to make for any disjunction all possibilities contradictory, and for any conjunction it suffices to make one conjunct contradictory, as and
-Of course, following a path in the graph, conjunctions for will become disjunctions for one of the confusing aspects of reasoning.
-(C1) and (C2) are antagonistic requirements.
-Satisfying (C1) creates new problems for (C2), and vice versa.
-In the limit, all requirements are satisfied, since all problems will solved in the next step, at the price of creating new problems, which will be satisfied in the next step.
-Note that we don’t approximate, as long we are finite, we always have only in the infinite case we have
-The right level of abstraction:
-The author tried first to work towards a representation result using single arrows.
-However, the following result suggests that one should work with (perhaps composite) paths instead of arrows: a single negative arrow may be replaced by the following diagram:
-So
-Thus, we may replace negative arrows recursively by arbitrarily deep constructions using negative arrows again, complicating the same construction arbitrarily.
-The same is true for all other logical operators.
-(See Section 4 in Sch23b.)
-4.2 In More Detail
-We now consider the inductive construction of the Yablo structure, see Diagram 4.1 (page 4.1) and Diagram 4.2 (page 4.2)
-In (2) and (5), we have unrelated points and we may consider the model sets to be orthogonal. - We only indicate model sets briefly, without going into details (U will stand for the universe).
-abbreviated
-has to lead to a contradiction for
-preparation,
-was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
-finish,
-(branch has to be contradicted for add
-has to lead to a contradiction for
-preparation,
-finish,
-(branch has to be contradicted for add
-(branch has to be contradicted for add
-Construction 4.1
-Summary:
-Thus, there are arrows for all and the construction is transitive for the
-Every is head of a Yablo cell with knee
-Thus, every arrow from any to any goes to the head of a Yablo cell, and not only the arrows from
-This property is “accidental”, and due to the fact that for any arrow there is also an arrow and property (2) holds for by prerequisite.
-The construction has infinite depth and branching.
-Diagram 4.1
-Diagram Inductive Construction 1
-Diagram 4.2
-Diagram Inductive Construction 2
-4.2.1 Composition
-The prerequisites for the constructions in the diagrams Diagram 4.1 (page 4.1) etc. are negative Paths, und formulas.
-These prerequisites hold also in our generalization to paths (modulo some simplifications for origins at etc.), so these constructions are also valid in the generalization to paths.
-Remark 4.1
-The condition “Existence of negative paths” is not trivial.
-Suppose that and are both negative, and is the only path from to then the condition is obviously false.
-The author does not know how to characterize graphs which satisfy the condition.
-Some kind of “richness” for the graph will probably have to hold.
-(This is covered in Yablo’s construction by transitivity.)
-Example 4.1
-There is a fundamental difference between the contradictions (for and
-Consider the case The path is blocked at So we cannot go to by two contradicting possibilities.
-The path is, however, not blocked at (or elsewhere), so we have two contradicting possibilities to go from to This basically behaves like and we have an escape possibility.
-So, if we try to append the same construction at we have an escape possibility, always making the origin of the construction negative.
-This is not true for We can append the same construction at as will always be at the opposite polarity of due to , the alternative path ends at so there is no escape.
-So a negative start at the original diagram leads to a positive start in the appended diagram.
-Diagram 4.3
-Diagram Postponed Contradictions
-4.3 Beyond Conjunctions
-We work with disjunctive normal forms.
 Consider
@@
 (In the case we have no independent parts, so the “problem” just discussed will not occur.)
-5 Conjecture
-Consider structures of the type (like ’s), with one small addition: we also allow formulas of the type either as (stand alone), or as one of several elements of a conjunction where is of the type
-(This is more general than adding the constant TRUE, as it permits chaining of the new formulas, but we also add TRUE for terminal
+1.2.3 The Graph Side
+We work with directed, acyclic graphs. They will usually have one root, often called In diagrams, the graphs may grow upward, downward, or sideways, we will say what is meant.
+Definition 1.1
+Nodes stand for propositional variables.
+If a node is not terminal, it has also a propositional formula attached to it, sometimes written with the meaning abbreviated The successors of are the propositional variables occuring in Thus, if and are the only successors of in may be but not
+Usually, the are (possibly infinite) conjunctions of propositional variables or (in most cases) their negations, which we write etc.
+We often indicate the negated variables in the graph with negated arrows, like etc.
+Thus, usually stands for more precisely.
+Example 1.2
+in above definition cannot be replaced by etc, as this example shows (we argue semantically, with the central conflictual construction in Yablo’s paper, Yablo Cell, see Section 1.2.1 (page 1.2.1), (2)):
+Let so so we have consistency, and it does not work for the construction.
+Example 1.3
+Consider the basic construction of a contradiction (used by Yablo and here, defined “officially” in Definition 2.1 (page 2.1).
+stands for so which is impossible.
+If we negate then so is possible.
+From the graph perspective, we have two paths in from to and
+We add now to so thus so and and is not contradictory any more.
+Definition 1.2
+(This applies only to unique occurrences of a variable in the formula attached to another variable.)
+We can attribute a value to a path expressing whether it changes a truth value from the beginning to the end. does not change the value of compared to that of does. We say or positive (negative) path.
+Formally:
+Let be paths as usual.
+If then if then
+Let be the concatenation of and Then iff and - otherwise.
+If all arrows are negative, then iff the length of is even.
+Definition 1.3
+We call two paths with common start and end contradictory, and the pair a contradictory cell iff The structures considered here will be built with contradictory cells.
+Remark 1.3
+Note that the fact that and are contradictory or not is independent of how we start, whether for both or for both
+We saw already in Example 1.3 (page 1.3) that it is not sufficient for a “real” contradiction to have two contradictory paths.
+We need
+(somewhere) an “AND”, so both have to be valid together, an “OR” is not sufficient,
+we must not have a branching with an “OR” on the way as in an “escape” path, unless this leads again to a contradiction.
+Fact 1.4
+(Simplified).
+Consider three paths, for simplicity with same origin, i.e.
+No contradiction loops of length 3.
+Suppose they meet at a common point, i.e. Then it is impossible that contradicts contradicts contradicts (at (“ contradicts ” means here that for some but one has value the other value -.)
+(Trivial, we have only 2 truth values).
+Suppose, first and meet, then (or and meet, but once they meet, they will continue the same way (e.g., if then for all ).
+Then it is again impossible that contradicts contradicts contradicts and continue to be the same but with different truth values until they meet so it is the same situation as above.)
+Above properties generalize to any loops of odd length (with more paths).
+This does not hold when the paths may branch again after meeting, as the next Example shows.
+Example 1.4
+(Example 7.4.2 in [Sch22].)
+Let
+then contradicts in the lower part, and in the upper part, contradicts and in the upper part, contradicts in the lower part.
+Obviously, this may be generalized to paths.
+Consider Yablo’s original construction:
+Example 1.5
+Let the nodes be and the arrows for expressed as a relation by and as a logical formula by
+Thus For any we have a contradiction by and for any and for any for a suitable
+Example 1.6
+This Example shows that infinitely many finitely branching points cannot always replace infinite branching - there is an infinite “procrastination branch” or “escape branch”.
+This modification of the Yablo structure has one acceptable valuation for
+Let as usual, and introduce new
+Let with
+If then by so, generally,
+if then and
+If then so if etc., so, generally,
+if then or
+Suppose now then for all and for all By there is and a contradiction, or for all again a contradiction.
+But is possible, by setting and for all
+Thus, replacing infinite branching by an infinite number of finite branching does not work for the Yablo construction, as we can always chose the “procrastinating” branch.
+Diagram 1.1
+Escape Path
+Remark 1.5
+For a representation result, we need enough paths in the graph considered to form contradictions.
+This is not trivial.
+Suppose that and are both negative, and is the only path from to then the condition is obviously false.
+The author does not know how to characterize graphs which satisfy the condition directly.
+Some kind of “richness” for the graph will probably have to hold.
+(This is covered in Yablo’s construction by transitivity.)
+1.2.4 Interplay of the Graph and the Logical Side
+We can either think on the logical level with formulas, or graphically with conflicting paths, as in the following Fact.
+We need infinite depth and width in our constructions:
+Fact 1.6
+The construction needs infinite depth,
+the logic as used in Yablo’s construction is not compact,
+the construction needs infinite width, i.e. the logic cannot be classical.
+Proof
+Let be a minimal element, then we can chose an arbitrary truth value and propagate this value upwards. If there are no infinite descending chains, we can do this for the whole construction.
+The logic as used in Yablo’s construction is not compact:
+Trivial. (Take This is obviously inconsistent, but no finite subset is.).
+It is impossible to construct a Yablo-like structure with classical logic:
+Take an acyclic graph, and interpret it as in Yablo’s construction.
+Wlog., we may assume the graph is connected.
+Suppose it shows that cannot be given a truth value.
+Then the set of formulas showing this does not have a model, so it is inconsistent.
+If the formulas were classical, it would have a finite, inconsistent subset, Define the depth of a formula as the shortest path from to this formula.
+There is a (finite) such that all formulas in have depth Give all formulas of depth (arbitrary) truth values, and work upwards using truth functions.
+As the graph is acyclic, this is possible.
+Finally, has a truth value.
+Thus, we need the infinite
+The following example illustrates (partially) unfinished contradictions, and resulting escape possibilities.
+Example 1.7
+There is a fundamental difference between the contradictions (for and
+Consider the case The path is blocked at So we cannot go to by two contradicting possibilities.
+The path is, however, not blocked at (or elsewhere), so we have two contradicting possibilities to go from to This basically behaves like and we have an escape possibility.
+So, if we try to append the same construction at we have an escape possibility, always making the origin of the construction negative.
+This is not true for We can append the same construction at as will always be at the opposite polarity of due to , the alternative path ends at so there is no escape.
+So a negative start at the original diagram leads to a positive start in the appended diagram.
+Diagram 1.2
+Diagram Postponed Contradictions
+1.2.5 A Third Truth Value
+See also Section 7.1.2, Fact 7.1.1 in [Sch22] for a less systematic approach.
+Remark 1.7
+We consider a third truth value, in addition to and
+means for a formula that neither nor are true (in the given set of formulas). Of course, this is impossible in a classical framework.
+Let be the truth value of
+Let be an abbreviation for and
+Let
+We have
+analogously.
+Likewise
+We thus have:
+and
+Using the operations for truth values
+and above results, we have
+(or and thus the order
+(and finally the inverse operations
+1.3 Strategy
+These comments will be clearer in hindsight, after having read the details. Still the best place to put them seems to be in general remarks, and go into details later on.
+The comments on “ is false ” show that there is no other way to express this in an acyclic graph of our type.
+The comments on the construction in the limit solve a seeming contradiction by chosing the right level of abstraction, i.e. by arguing on the level of instances.
+1.3.1 Expressing “x is false” In an Acyclic Graph
+In an acyclic graph, we cannot express something like we have to use different means to express that is impossible.
+As we can always fill in truth values (including from the bottom, we cannot force by looking upwards of (or sideways, i.e. there is no monotone path connecting with some So we have to look downward.
+We may use or if we have the constant We will see later, that this not interesting, see Definition 2.1 (page 2.1) and Remark 2.1 (page 2.1).
+This is a contradiction made of (graphically speaking) two paths of length one each.
+Interesting contradictions are made of slightly more complicated paths, like the Yablo Cell meaning
+Basically, the only way to express that is impossible, is by the logical formula
+A slightly different argument: How can we express that is wrong in a finite acyclic graph?
+Consider the set of minimal elements below , and the formula corresponding to the graph in DNF of these elements.
+This might be
+(the graph is trivial, just
+or a tautology or contradiction.
+a disjunction of conjunctions of elements in , or their negations. By acyclicity,
+This argument is not totally satisfactory. For this reason, we reformulate:
+We examine graphs, where there existe a node such that as well as -x will lead to a contradiction.
+1.3.2 Satisfying the Antagonistic Requirements (C1) and (C2) In the Limit
+We want to make and contradictory.
+We know that this is impossible in finite graphs, so we have to show how we solve the problem in the (infinite) limit.
+We proceed inductively as Yablo did, but we also see that we basically have no other choice (modulo minor modifications).
+So we want to express by an infinite graph that both and are inconsistent.
+As we saw in Section 1.3.1 (page 1.3.1), this must be expressed by suitable contradictions.
+Therefore, we treat such contradictions as basic building blocks, as Yablo did.
+The problems are then
+- (1)
+How do we combine them to obtain a suitable graph expressing that both and are impossible?
+The problem is that, if is contradictory, then is a tautology in classical logic, and thus in the theory expressed by a finite acyclic graph.
+So we alternate between and a tautology, and it is not clear what such constructions are in the limit, when we leave classical logic.
+This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 3 (page 3).
+A summary: we have to look at instances.
+- (2.1)
+- (2.2)
+But is now a branching point. The branch continues to be satisfied for (C1), but the (new) branch opens a new problem for (C1), as is an OR for
+- (2.4)
+So we have new instances of problems for (C1), (C2), we solve the new instances, say for (C1), the old instances for both (C1) and (C2) stay solved.
+Thus, in the limit, new problems (instances of problems, more precisely) are created, they are solved in one of the next steps, problems once solved stay solved, so there remain no unsolved problems.
+The set of problems increases, but all are solved in the limit, and and are contradictory.
+We have (essentially) an alternation of creating and treating problems for (C1) and (C2): We cannot treat e.g. a (C1) problem, bevor it was created by solving a (C2) problem - and conversely.
+(Depending on organization, we have some liberty: e.g.
+solving (C1) creates a (C2) problem, but this needs TWO branches.
+So we might first treat these two branches ((C1) problems) before going back to the new (C2) problem created by the first new branch.)
+1.4 The “Right” Level of Abstraction
+1.4.1 Introduction
+The basic elements in Yablo’s construction are negative arrows, from which Yablo Cells are built.
+We show here that we can build arbitrarily complex structures equivalent to negative arrows, and as a matter of fact to any propositional logical operator.
+This suggests that the right level of abstraction to consider more general Yablo-like structures is not the level of single arrows, but rather of paths.
+Remark 1.8
+In the diagrams in Diagram 1.3 (page 1.3) the inner path is barred by , and the outer path contains 3 or 2 negations.
+Definition 1.4
+We define negations and their type.
+A negation (diagram) is a diagram all of whose arrows are of the kind.
+We define negation diagrams, or, simply, negations, and their type.
+An arrow is a negation of type 0.
+all negations inside are of type and at least one negation inside is of type
+Remark 1.9
+Thus, it seems very difficult to describe Yablo type diagrams on the level of single arrows. It seems to be the wrong level of abstraction.
+We now give some examples.
+1.4.2 Some Examples
+Example 1.8
+Negation (1)
+so and
+Negation (2)
+so and continue as above.
+Identity
+so and
+TRUE
+Thus, and
+so and
+Diagram 1.3
+Propositional Formulas 1
+Diagram 1.4
+Propositional Formulas 2
+1.5 Prerequisites
+We work with pure conjunctions of positive or negative propositional variables, like except for the following:
+we may use a constant TRUE or
+we may use formulas of the type
+so the following are also permitted: etc.
+This is used only to “neutralize” unwanted parts of the graph we characterize, see Fact 4.4 (page 4.4).
+- (2)
+- (3)
+- (4)
+2 Elementary Cells
+2.1 Introduction
+Elementary cells have to be simple constructions providing contradictions (in some cases, e.g.
+when the origin, usually called or so, cannot have a certain truth value, as this will lead, within the cell, to a contradiction).
+Thus
+They must be contradictory, and thus exclude at least one instance.
+In addition, they should be simple, e.g. they will not contain a sequence or so without branching, as we could replace this with
+(We will see later how to relinquish this requirement.)
+More importantly, we must be able to construct a Yablo-like structure, built up by this type of elementary cell only (and perhaps some simpler auxiliary structures).
+They must serve as building blocks for a suitable stucture.
+This prevents “cheating”, letting other constructions do the crucial work.
+- (4)
+2.2 Basic Discussion
+See Section 7.3 in [Sch22] for more material.
+We try to describe here the basic constructions of contradictions.
+As said in Fact 1.1 (page 1.1) (and re-written graph-theoretically), these consist of two branches with common origin, which meet again, and have different polarity. We call such constructions cells.
+First, we want to exclude some trivialities. We describe only the part beginning at the branching point, not before, not after.
+Second, we will define a simple hierarchy of such cells, and will allow a cell only the use of simpler cells as substructure.
+This prevents “cheating”.
+(See Section 2.1 (page 2.1).) Thus, we look at those cells only which allow to construct a Yablo-like structure without the use of more complicate cells.
+We first use (almost) only negative arrows, and nodes whose formulas are conjunctions. We will see how to generalize to more complicated paths.
+Definition 2.1
+We will argue that it suffices to consider these types of cells, see Remark 2.3 (page 2.3) and Section 2.4 (page 2.4).
+Furthermore, we can also neglect all but the (slightly generalized) Yablo Cells, as we will see in the present section.
+It will become clear in a moment that above cells are fundamentally different, but for this we have to consider the second requirement (C2).
+Remark 2.1
+An escape problem with for the trivial contradiction, so (See Definition 2.1 (page 2.1).) If we continue with so etc. we have an escape sequence etc., which never meets a contradiction.
+Thus, is not a contradiction cell for our purposes, has a model defined by the escape sequence.
+Suppose we have an escape problem, like thus
+Can we append a new finite structure to which reduces the possibilities to just one value, say for So, whatever the input, or the outcome is
+This, however, is not possible.
+There are, whatever the inner structure, just four possibilities: it might be equivalent to The first two cases are trivial. Suppose Then Suppose Then
+In general, suppose we have a partial graph with origin and end point (among perhaps other end points), and, say results in choices or (due to internal choices by OR).
+Example 2.1
+We consider now some simple, contradictory cells.
+They should not only be contradictory for the case but also be a potential start for the case without using more complex cells.
+(Otherwise, we postpone the solution, and may forget the overly simple start.)
+We will consider in Section 2.5 (page 2.5) a more complicated contradiction cell (“Diamond”), and discuss here only simple cells.
+Note that the following considerations apply also to cells formed by paths consisting of more than one arrow, and not only to cells formed by single arrows.
+See Section 2.4 (page 2.4).
+In particular, also in more complicated cells all sides have to negative.
+The cell with 2 arrows.
+It corresponds to the formula graphically, it has a positive and a negative arrow from to so exactly one of and is negative.
+If is positive, we have a contradiction.
+Of course, appending at a Yablo Cell (see below, case (2.3)) may be the beginning of a contradictory structure, but this is “cheating”, we use a more complex cell.
+Cells with 3 arrows.
+In the graph with not annotated arrows, they have the form -z, -z, where the lines may stand for or
+Again, we want a contradiction for positive, so we need an at For the same reason, the number of negative arrows may not be even. So assume the number of negative arrows is one or three.
+Three negative arrows:
+This is the original type of contradiction in Yablo’s construction
+This will be discussed in detail in the rest of the paper.
+But we see already that both paths, and change sign, so and will be positive, appending the same type of cell at and solves the problem (locally), and offers no escape.
+We call sometimes the head of the Yablo cell, its foot, and its knee.
+If we combine Yablo Cells, the knee for one cell may become the head for another Cell, etc.
+One negative arrow:
+We will show that all cases have an escape possibility, along the positive arrow (or arrows) originating in So they are not suitable for a construction using only this type of cell.
+Consider the case
+Upper part:
+Lower part:
+Consider the case
+Consider the case
+Here,
+Diagram 2.1
+Simple Contradictions
+Diagram 2.2
+Diagram ESC-1
+Diagram 2.3
+Diagram ESC-2
+Diagram 2.4
+Diagram ESC-3
+Remark 2.2
+The distinction between and i.e.
+between knee and foot, is conceptually important.
+In the case we have at a complete contradiction, at we have not yet constructed a contradiction.
+Thus, if we have at again an (as at this becomes an and we have to construct a contradiction for all (or not only for Otherwise, we have an escape possibility for Obviously, the contradiction need not be immediate at the important property is that ALL paths through all lead to a contradiction, and the simplest way is to have the contradiction immediately at - as in Yablo’s construction, and our saw blades.
+See also the discussion of “Saw Blades” in [Sch22], Section 3.
+As we work for a contradiction in the case x-, too, the simplest way to achieve this is to have a negative arrow and at again an This gives a chance to construct a contradiction at Of course, we have to construct a contradiction at too, as in the case we have an at
+Of course, we may have branches originating at which all lead to contradictions in the case so (resp. at is possible, too.
+But, for the simple construction, we need between and and at And this leads to the construction of contradictions for all (or as just mentioned above.
+2.3 Values for y and z in Yablo Cells
+We consider here what happens when we give and values In addition, we consider the variants and
+Remark 2.3
+False, True
+We present a systematic treatment of variants of the “environment” of the Yablo triangle (and the diamond).
+In the Yablo construction, we attach to the Yablo triangle to and constructions, which make We also consider here the cases where we end by or and set e.g. etc., see (3.1) below.
+Note that will always hold.
+We consider the following requirements:
+Do we have a contradiction for i.e. will not be true in the Yablo Cell?
+Will both branches lead to possible contradictions for
+Do we obtain
+- (2.4)
+We consider the cases in all possible combinations:
+Case is equivalent to case
+Thus, we look at e.g. case i.e. and etc., and consider whether the requirements in (2) above are satisfied.
+We first look only at the minimal requirements
+has to be contradictory, requirement (2.1) above.
+must not be classical (i.e. not or requirement (2.3) above.
+The cases:
+(2) (requirement (2.3))
+Consider so for any so these cases are classical.
+Consider For is classical, so is, too.
+so
+Thus, is the only non-classical case among
+Consider In cases so too.
+In case so
+In case so
+So is the only classical case among the
+So the interesting cases left are
+(1) (requirement (2.1))
+We now check for local contradiction in the case i.e. must not be
+may only be iff But if then only in case
+Next, we have to look at and see if every path from leads to a contradiction, requirement (2.2) above. So neither nor must be
+So and will not work.
+Consider i.e.
+Then in and so this does not work.
+so this works.
+so this works.
+Consider i.e.
+Then or and all cases work, except
+In summary: in exactly the following cases: so they are excluded.
+We summarize the results for the three requirements above:
+(2.3): are ok.
+(2.2): are not ok, so we learn nothing new beyond requirement (2.3).
+(2.1): Only is eliminated, so we learn nothing new beyond requirement (2.3).
+In summary, only satisfy all three requirements, (2.1), (2.2), (2.3).
+We look at these cases.
+Here, and in (3.6), stands for which is of type likewise as we will append in the full structure at and a construction with value
+In all cases
+Thus,
+Thus,
+Thus,
+identical to
+Thus,
+Thus,
+We look at the Diamond.
+Thus,
+The modified Diamond is the same, only so and so we have a Yablo cell.
+Summary of the (in this context) important constructions
+2.4 Yablo Cells Formed By Paths
+We consider a triangle , , where etc. may be composite paths. We know from Section 2.2 (page 2.2) that the paths have to be negative, but they may be composed of positive and negative arrows.
+We require that is of type and that the triangle is contradictory for the case as well as for both paths und have to lead to a contradiction (conditions (C1) and (C2)).
+Note that and have to be the first nodes on above paths to be contradictory for - otherwise we will not reach (or the path is barred before.
+2.4.1 Branching points
+We have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).
+Branching at
+Suppose we branch at so we have, in addition to the triangle, some
+We know that is negative, so, if then This is the situation of the Yablo construction, see Section 3 (page 3). So just a short comment here.
+Thus, we also need for a contradiction, this leads to infinite branching at
+A new path as in Yablo’s construction leads to infinite branching and non-classical logic.
+We may branch on the already existing oder , and continue to build a (finite) contradiction to
+This however, is possible only a finite number of times, e.g. then etc., see Example 1.6 (page 1.6). This constructs an infinite sequence of choices which offers an escape possibility.
+Thus we need infinite branching (and non-classical logic).
+We see that we may consider only the new paths they form an infinite subsequence of the construction.
+By full transitivity of the final construction, any infinite subsequence of the construction has the same properties as the full construction.
+2.4.2 Paths
+Additional branching on , e.g. , ,
+Case 1: OR
+This generates a copy of the structure for
+We construct for as for , i.e.
+Case 2: AND
+So OR, so we make a copy of the structure for consider e.g. , ,
+Note: We simplify, suppose e.g.
+that we have then we do not need another contradiction at the one at suffices (condition (C1)).
+We need, however, that all branches lead to contradictions (condition (C2)).
+We branch on , and have , .
+Case 1: OR:
+We make a copy of the structure for
+Case 2: AND:
+So x- OR:
+We make a copy of the structure for
+Consider e.g. , , .
+As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (” pipeline “), we may assume for simplicity that we always start contradictions at
+Diagram 2.5
+2.5 The Problem with Diamonds
+Remark 2.4
+We have a conflict between the diamonds starting at and and the diamond starting at
+If then y- and As the choices at and are independent, any branch -z, combined with any branch -z, must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
+However, if is positive, the diamond -z, has to be contradictory, so not both branches may be negative.
+A solution is to “synchronise” the choices at and which can be done e.g. by the formula
+and This is a different type of formula, moreoever corresponding arrows from are missing.
+Even if we do not consider the paths for the diamond starting at we have and so we can make and false, and chose and which results in an escape possibility (we have to consider new diamonds starting at etc.)
+This formula has an additional flaw, we want to speak about paths we chose, and not just their end points.
+This can be done with introducing additional points on the arrows, e.g.
+the arrow is replaced by now we can speak about the arrow we have given it a name.
+We may also put such information in a background theory, which need no be negated.
+But all this leads too far from the basically simple formalism of Yablo’s construction, so we will not discuss this any further.
+(Yablo’s construction does not need synchronisation, this is done automatically by the universal quantifier.)
+The simplified version with recursion on one side only, here on the left only, is nothing but the Yablo triangle:
+Thus, we conclude that the original version with 4 points and full recursion is beyond our scope, and the simplified version (see above) is nothing but the Yablo triangle.
+Diagram 2.6
+Nested Diamonds, Details
+Diagram 2.7
+Diamond Essentials, Recursion on the Left, Simple Contradiction on the Right
+3 Combining Cells
+3.1 Overview
+We follow Yablo’s construction.
+(C1): We have to make contradictory.
+We first try to make “directly” contradictory. That is, we will not first go to and make contradictory, nor will we make a disjunction at and then make each disjunct contradictory.
+(This is not an important simplification, as more complicated constructions will finally use our simple one, too. See e.g. Diagram 1.2 (page 1.2) and its discussion.)
+As Yablo did, we set and
+The first step is done, is contradictory. Probabilistically, of the possibilities via are contradictory.
+(C2) We have to make contradictory.
+Now, must be contradictory, too. As we do not know which one is positive, we must make both and contradictory - just as a disjunction is false if every disjunct is false.
+Consider, e.g., Again, we will make directly contradictory (as above for so is a conjunction, say (we might re-use here, but this will not be important) and add e.g.
+and to for the contradiction, so we have
+But now, we have destroyed the contradiction at as for is a disjunction, and have to make ALL possibilities for contradictory for so far only the possibility is contradictory.
+Probabilistically, of the possibilities at are open.
+To create a contradiction for the possibility we introduce new contradictions by adding to so
+Etc.
+See below and Section 3.3.3. in [Sch23b].
+Note that for (C1), i.e. for and for (C2), i.e. for we have to make for any disjunction all possibilities contradictory, and for any conjunction it suffices to make one conjunct contradictory, as and
+Of course, following a path in the graph, conjunctions for will become disjunctions for one of the confusing aspects of reasoning.
+(C1) and (C2) are antagonistic requirements.
+Satisfying (C1) creates new problems for (C2), and vice versa.
+In the limit, all requirements are satisfied, since all problems will solved in the next step, at the price of creating new problems, which will be satisfied in the next step.
+Note that we don’t approximate, as long we are finite, we always have only in the infinite case we have
+3.2 In More Detail
+We now consider the inductive construction of the Yablo structure, see Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2)
+In (2) and (5), we have unrelated points and we may consider the model sets to be orthogonal. - We only indicate model sets briefly, without going into details (U will stand for the universe).
+abbreviated
+has to lead to a contradiction for
+preparation,
+was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
+finish,
+(branch has to be contradicted for add
+has to lead to a contradiction for
+preparation,
+finish,
+(branch has to be contradicted for add
+(branch has to be contradicted for add
+Construction 3.1
+Summary:
+Thus, there are arrows for all and the construction is transitive for the
+Every is head of a Yablo cell with knee
+Thus, every arrow from any to any goes to the head of a Yablo cell, and not only the arrows from
+This property is “accidental”, and due to the fact that for any arrow there is also an arrow and property (2) holds for by prerequisite.
+The construction has infinite depth and branching.
+Remark 3.1
+Our construction is somewhat arbitrary. We construct a contradiction for by and use the arrow as part of the contradiction for This, of course, is a special case.
+Consider Diagram 3.3 (page 3.3). Here, the contradiction for is as above, but we begin a totally new contradiction for likewise for etc.
+But, we see, the old construction of Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2) is part of the more general construction of Diagram 3.3 (page 3.3), so the old construction is really minimal.
+We have to be a bit careful here. If we set etc. then we have etc. Instead, we may set so we have as desired, or we append at recursively the same construction - as in “Saw Blades”.
+Diagram 3.1
+Diagram Inductive Construction 1
+Diagram 3.2
+Diagram Inductive Construction 2
+Diagram 3.3
+Diagram Inductive Construction Core
+4 Result
+Consider structures of the type (like ’s), with one small addition: we also allow formulas of the type either as (stand alone), or as one of several elements of a conjunction where is of the type It is important that these additions also cooperate with our truth value
+(This is more general than adding the constant TRUE, as it permits chaining of the new formulas, but we also add TRUE for terminal nodes.)
 In addition, we assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it.
@@
 See here e.g.
-the beginning of Section 3.1 (page 3.1), and Remark 3.1 (page 3.1), Definition 2.4 (page 2.4), and Example 4.1 (page 4.1), with Diagram 4.3 (page 4.3).
+the beginning of Section 2.2 (page 2.2), and Remark 2.1 (page 2.1), and Example 1.7 (page 1.7), with Diagram 1.2 (page 1.2).
 Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
-5.1 Conjecture: A Suitable Injection Does the Job of Characterisation
+4.1 A Suitable Injection Does the Job of Characterisation
 Consider Yablo’s original structure YS, and some directed acyclic graph YS’.
-Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory iff
+Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory iff the following conditions hold.
+Condition 4.1
 There is an injection from YS to YS’, more precisely from the nodes of YS to the nodes of YS’, s.t., if there is a path from to in YS, then there is a path from to
 For THIS injection, there is a valuation of the arrows in (the other arrows are not interesting) s.t., if form a contradiction in YS, then likewise for (Contradictions are preserved under
-5.2 Idea of Proof (Outline)
-Fact 5.2
+4.2 Proof
+Fact 4.2
 Yablo’s structure is the simplest structure showing that both and are contradictory.
@@
 Earlier sections of this paper. More precisely:
-- (1.1)
-- (1.2)
-- (1.3)
+- (1)
+- (2)
+- (3)
 Remark 4.3
@@
 EG: so , this works well classically, but also for as
-We may also continue beyond hinaus: or then too.
+We may also continue beyond : or then too.
 If we have mixed chains of and the last element will decide the result (trivial).
-Fact 5.4
-Adding new formulas inside a Yablo Triangle does not improve anything, it only destroys.
-It leads either to no contradiction, or to a contradiction of the simplest type, and thus to an escape situation, see Definition 2.4 (page 2.4).
-There are no other new contradictory pairs of paths.
-Proof
-Consider a YC
-Insert in ab, or and So, if then if then
-Consider now insertions of xy, or xz, yz (analogously) and arbitrary combinations thereof.
-xz: If then of then so we have a trivial contradiction and an escape possibility.
-xy: xy as so If then if then we have again a trivial contradiction and an escape possibility.
-yz: so If then if then we have again a trivial contradiction and an escape possibility.
-Combinations:
-For combinations of xy and yz, see (2.5) above.
-xy and xz: the result is TRUE (both positive) or trivial contradiction with ESC.
-xz and yz: This is analogous to xy and xz.
-Let be such paths originating in containing some xy as in Fact 5.4 (page 5.4).
-Then the same considerations as above apply, so we have either no contradiction, or the trivial one with an escape possibility.
-Similar results hold when using the constant TRUE: If has -TRUE in one branch, then will have TRUE in this branch, violating (C2).
-If has TRUE in one branch, then this branch is irrelevant, and we have no contradiction.
-So TRUE must not occur.
-Extend the interpretation beyond
-Let If has no successor, then set
-If has successors etc., then set
-If then has successors. Take the conjunction of the following: if if then
-We have to show that still does what it should.
-During construction, will be and after completion, And any will be TRUE or By the combination with AND, will not change.
+Fact 4.4
+Consider (without the other elements etc. in
+Now consider the full YS’, and attribute values to the nodes as follows, respecting the values already given in (1):
+If is minimal, then
+Otherwise:
+If dann
+arrow
+Note that these arrows have no value, i.e. we have no distinction between and
+If then
+: arrow
+: arrow
+: arrow
+Note that arrows inside have a value, these values have to be preserved. The other arrows have no value, and we treat them as positive.
+Moreover, we did not distinguish between minimal and not minimal this does not matter. If we want to distinguish we have to differentiate the last conjunct further.
+- (3)
+Remark 4.5
+Asssume for simplicity that (arrows may be negative, too), likewise for etc.
+We have paths etc., assume for simplicity that the only new intermediate node is respectively).
+Let is not in
+We set arbitrary, Thus, so and so
+Of course, instead of going through the additional path might also go through etc. This changes nothing. We have neutralised the path not in
+Diagram 4.1
+Diagram
+Diagram 4.2
+Diagram

### v5

@@
 Abstract
-We continue our analysis of Yablo’s coding of the liar paradox by infinite acyclic graphs, [Yab82].
-The present notes are based on and continue the author’s previous results on the problem.
-In particular, our approach is often more systematic than before.
+We show that the existence of a node with attached formula in an acyclic graph such that neither nor its negation are consistent is equivalent to the existence of subgraph equivalent to Yablo’s construction.
 Contents
 1 Introduction
-Please note that arXiv pdf files sometimes have problems with some of the diagrams. In that case, please consult the html version.
 1.1 Outline
 These notes are based on [Sch22] and [Sch23b], the reader may have copies of both ready for more examples and discussions.
-See also the author’s earlier versions on arXiv.
-In particular, the reader is referred there for a discussion of more complicated formulas attached to the nodes.
-We liberally use material from these sources.
+See also [Sch25].
+(Please note that arXiv pdf files sometimes have problems with some of the diagrams.
+In that case, please consult the html version.) We liberally use material from these sources.
 All these papers are, of course, based on Yablo’s seminal [Yab82].
+Prerequisites:
+All graphs considered will by acyclic.
+- (1.2)
+For ease of exposition, we introduce a third truth value, as abbreviation for “ and are inconsistent”.
+We will use only one type of contradictory cell in a graph. This prevents “cheating”.
+Contradictions in a graph will begin immediately, and not after some arbitrary path. This simplifies constructions, and is not an important restriction.
+- (1.6)
 In Section 1.2 (page 1.2), we first present the background, its logical and graph theoretical side, and their interplay.
 We also discuss a useful third truth value, which expresses that neither nor is consistent.
-In most cases, we will work with simple conjunctions (with and without negations).
+In Section 1.3 (page 1.3), we discuss our strategy, and discuss the antagonistic conditions (C1) and (C2) and how to satisfy them in an inductive construction.
+(We need infinite depth and width in the construction, see e.g.
+[Sch25], Fact 2.4, or [Sch23b], Fact 1.4).
 In Section 2 (page 2), we discuss elementary contradictions (cells).
 Essentially, these are variations of Yablo’s triangles.
-The simplest contradictions are too simple, they have an “escape possibility”, i.e., combining them will leave one possibility without contradictions.
+The trivial contradiction, see Definition 2.1 (page 2.1), is too simple, it has an “escape possibility”, i.e., combining them will leave one possibility without contradictions.
+See Definition 1.6 (page 1.6).
 Diamonds, constructions with four sides, will not work because they need a “synchronization”, which is not available in our framework.
-In Section 3 (page 3), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox. Basically, we follow Yablo’s construction.
+See Section 2.5 (page 2.5).
+We also show that all sides of the triangles have to be negative, see Remark 2.2 (page 2.2).
+Section 2.3 (page 2.3) discusses further details of these triangles.
+We generalize to triangles formed by paths in Section 2.4 (page 2.4).
+In Section 3 (page 3), we discuss how to put contradiction cells together to obtain graphs, which code the liar paradox.
+Basically, we follow Yablo’s construction.
 We then generalize slightly, see Remark 3.1 (page 3.1), but show that this general construction contains Yablo’s construction, so it is minimal.
-(See also the discussion of “Saw Blades” in Section 7.6 of [Sch22] and Section 3 in [Sch23b].)
-In both cases, we show that - within the chosen variant - the construction is necassary.
-- (4)
-There is an important property for the graphs, which we need, probably a sort of “richness”, see Remark 1.5 (page 1.5).
-This property holds trivially in Yablo’s construction by transitivity.
-In our approach, it will hold by the construction of a suitable injection, see Section 4 (page 4), Condition 4.1 (page 4.1) (2).
+This is perhaps the central result of the paper.
+This fact is essentially re-used in Section 4 (page 4).
+(See also the discussion of “Saw Blades” in Section 7.6 of [Sch22] and Section 3 in [Sch23b].) In both cases, we show that the construction is necessary.
+In Section 4 (page 4), we show that
+a graph without valuation can be given a valuation such that there is an element such that neither nor are consistent, if there is a suitable injection from Yablo’s structure into and a partial valuation for the nodes of the paths of the injection, such that contradictions are preserved.
+We complete the partial injection by a trivial extension.
+conversely, if there is a graph with (complete) valuation with some node such that neither nor are consistent, then we can construct a subgraph which corresponds again to an injection of the Yablo structure.
+We use here again the minimality of Yablo’s construction, this time imside the given graph.
 1.2 Basic Definitions and Facts
-1.2.1 Terminology
-Yablo’s construction
-Yablo’s construction is an infinite directed graph with nodes and (negative) arrows for
-Nodes stand for propositional variables, and the meaning of arrows is that variable occurs (negatively) in the formula attached to in most cases it expresses (We will identify for simplicity nodes with their attached variables.)
-It is easy to see that no node may be true, nor false: Suppose is true, than all are false, so must be false, then contradiction, as all must be false.
-Suppose is false, then some must be true, again a contradiction (to the above).
-Yablo Cell, Yablo Triangle
-The basic contradiction structure has the form with the meaning (so and
-We sometimes call the head, the knee, and the foot of the Yablo Cell (or Triangle).
-Conditions (C1) and (C2)
-We sometimes write when we want to emphasize that we assume is positive, similarly for the negative case.
-Consider in the Yablo construction (but this applies to all
-We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction.
-(More precisely, see Section 3.1 (page 3.1), (3).)
-1.2.2 The Logical Side
-On the logics side, we work with propositional formulas, which may, however, be infinite.
-We will see that we need infinite formulas (of infinite depth) to have nodes to which we cannot attribute truth values. See Fact 1.6 (page 1.6).
 Notation 1.1
-When we give nodes a truth value, we will use (and etc. if has the form to denote the case and etc. for the case
-We will work here with disjunctive normal forms, i.e.
-with formulas of the type where and the are propositional variables or negations thereof - most of the time pure conjunctions of negations of propositional variables.
-Fact 1.1
-Let where the are propositional variables or negations thereof.
+We consider directed, acyclic graphs. They will usually have one root, often called Nodes and arrows may be labelled.
+If the nodes are labelled, we identify nodes with their labels, unless said otherwise.
+The labels of nodes stand for propositional variables or constants The propositional variables will be unique in a given graph.
+Moreoever, to is associated a formula the direct successors of a node determine the propositional variables or constants occurring in The theory of a graph is the union of its a node in the graph.
+We sometimes write for the theory of a graph
+We sometimes write if occurs negatively in
+In many cases is a conjunction of positive or negative occurrences of the direct successors of If not, we will say so expressly. For instance, the graph will express that we often write instead of
+Definition 1.1
+The pure graph, i.e.
+just nodes and arrows, tells only which variables occur in the associated formulas We call the annotations or valuations, the pure graph graph without valuations or annotations, and the full graph annotated graph or graph with valuations, etc.
+If is the pure graph, will sometimes be an annotated version of and conversely.
+Definition 1.2
+Let be a node in a fixed graph as above.
+We say that is contradictory (relative to iff the assumption leads to a contradiction in Likewise for is contradictory, iff the assumption is contradictory in this theory.
+Truth values:
+We set iff is contradictory, and iff not, likewise for
+We introduce a third truth value which is an abbreviation for and - relative to of course. See Remark 1.2 (page 1.2).
+When the context is clear, we sometimes simplify to etc.
+Contradictions via paths - we will change between the graph and the logical formulation, whatever seems intuitively better:
+Such contradictions have in a graph the following form: there is a node and a node and two paths from to which branch at and join again at and they contradict each other, i.e.
+one branch results in the other in
+These are the basic contradictions.
+- (3.2)
+Consider again Diagram 3.1 (page 3.1), part (2.1).
+The whole diagram is not contradictory, as we opened a new possibility at which becomes impossible only in part (3) of this diagram.
+So, as a detail, the paths are contradictory, but the possibility is closed only in (3), so the contradiction of the whole diagram is only complete in (3).
+If you prefer to think probabilistically, in part (2.1) of the cases are contradictory etc.
+See also Section 1.3.1 (page 1.3.1) for opening and closing new possibilities.
+Definition 1.3
+(Yablo, see [Yab82].)
+We call the following graph the Yablo structure, YS:
+Written as graph: We use negated arrows, and positive arrows, in graphs as this improves readability considerably (see Notation 1.1 (page 1.1).)
+Written as set of propositional formulas:
+Contradictions in YS:
+Suppose is a node in the YS. Then and which is contradictory.
+But starting with will also result in a contradiction:
+suppose is true, then is contradictory, as we just saw, so is contradictory, too.
+Note that we abused notation here, where may be an arrow in a graph, or an implication in a logical formula. Context will tell.
+The contradictions have the form this is the basic contradiction “cell”, we call such cells Yablo cells, ’s, or Yablo triangles.
+We sometimes call the head, the knee, the foot of such ’s. This is relative to the cell discussed. When we combine such cells, the knee or foot of one cell may become the head of the next cell, etc.
+Remark 1.1
+Let where the are propositional variables or negations thereof, thus, is in disjunctive normal form, DNF.
 Let - the set of choice functions in the sets where
@@
 (Argue semantically with the sets of models and general distributivity and complements of sets.)
-Contradictions will be between two formulas only, one a propositional variable, the other the negation of the former.
-Fact 1.2
-Let and we want to show that is contradictory, then it suffices to show that a subset of the is inconsistent - this corresponds to the fact that (X any truth value.)
-Let and we want to show that is contradictory, then we have to show that ALL are contradictory. Again, this corresponds to
-The following example illustrates the propagation of required contradictions.
-Example 1.1
-Consider
-As has to be contradictory, both components and have to be contradictory.
-As has to be contradictory, all components, etc., have to be contradictory.
-For this, it is sufficient and necessary that in each component, one element is contradictory.
-This is equivalent to the condition that (both and or (both and are contradictory.
-Thus
-For both
-and
-have to be contradictory, this is done as usual by
-and
-For we have to attach a contradiction at either
-(-1) and
-or
-(-2) and
-We thus have to examine the combinations
-(-1)
-(-2)
-(-1)
-(-2)
-The combinations (-1)) and (-2)) are as usual, the other ones are new.
-It seems that there might be a problem.
-In the cases and seems to be independent from and both solvable by classical logic, a contradiction.
-But all cases are parts of and and (-1)) and (-2)) need simultaneous treatment of the and the - part.
-So the confusion is only transient.
-(In the case we have no independent parts, so the “problem” just discussed will not occur.)
-1.2.3 The Graph Side
-We work with directed, acyclic graphs. They will usually have one root, often called In diagrams, the graphs may grow upward, downward, or sideways, we will say what is meant.
-Definition 1.1
-Nodes stand for propositional variables.
-If a node is not terminal, it has also a propositional formula attached to it, sometimes written with the meaning abbreviated The successors of are the propositional variables occuring in Thus, if and are the only successors of in may be but not
-Usually, the are (possibly infinite) conjunctions of propositional variables or (in most cases) their negations, which we write etc.
-We often indicate the negated variables in the graph with negated arrows, like etc.
-Thus, usually stands for more precisely.
-Example 1.2
-in above definition cannot be replaced by etc, as this example shows (we argue semantically, with the central conflictual construction in Yablo’s paper, Yablo Cell, see Section 1.2.1 (page 1.2.1), (2)):
-Let so so we have consistency, and it does not work for the construction.
-Example 1.3
-Consider the basic construction of a contradiction (used by Yablo and here, defined “officially” in Definition 2.1 (page 2.1).
-stands for so which is impossible.
-If we negate then so is possible.
-From the graph perspective, we have two paths in from to and
-We add now to so thus so and and is not contradictory any more.
+Example:
+Consider the Yablo cell with each formula associated with of the type We then have 16 possible combinations like etc., 7 are in 9 in is in the 9 combinations of and the 9 combinations of By is in the 7 positive combinations of so we have again a contradiction, as usual.
+(Alternatively, we may simplify, and work instead of with many combinations, e.g. just with the totally positive combination for and with the totally negative combination
+The generalization to arbitrary DNF’s is trivial.
+Thus, working with DNF’s changes nothing compared to the Yablo construction, working with All we need is a suitable negation.
+Remark 1.2
+The truth value has the following properties:
+(and, of course
+the following order is compatibel with and i.e. for any truth values and
+This completely describes the behaviour of
+This is the only 3-valued logic with the property and compatibility with and It is thus identical to the Strong Kleene logic Ks, as was pointed out by a referee.
+Proof:
+For more details, see [Sch25]
+trivial
+Consider two copies of ’s, YS and YS’, beginning with and respectively, and the following two structures (a) with (b) with
+Consider a new origin, and arrows from to as follows: and and
+Details are left to the reader.
+Easy.
 Definition 1.4
@@
 Let be the concatenation of and Then iff and - otherwise.
-If all arrows are negative, then iff the length of is even.
-Definition 1.3
-We call two paths with common start and end contradictory, and the pair a contradictory cell iff The structures considered here will be built with contradictory cells.
-Remark 1.3
-Note that the fact that and are contradictory or not is independent of how we start, whether for both or for both
-We saw already in Example 1.3 (page 1.3) that it is not sufficient for a “real” contradiction to have two contradictory paths.
-We need
-(somewhere) an “AND”, so both have to be valid together, an “OR” is not sufficient,
-we must not have a branching with an “OR” on the way as in an “escape” path, unless this leads again to a contradiction.
-Fact 1.4
-(Simplified).
-Consider three paths, for simplicity with same origin, i.e.
-No contradiction loops of length 3.
-Suppose they meet at a common point, i.e. Then it is impossible that contradicts contradicts contradicts (at (“ contradicts ” means here that for some but one has value the other value -.)
-(Trivial, we have only 2 truth values).
-Suppose, first and meet, then (or and meet, but once they meet, they will continue the same way (e.g., if then for all ).
-Then it is again impossible that contradicts contradicts contradicts and continue to be the same but with different truth values until they meet so it is the same situation as above.)
-Above properties generalize to any loops of odd length (with more paths).
-This does not hold when the paths may branch again after meeting, as the next Example shows.
-Example 1.4
-(Example 7.4.2 in [Sch22].)
-Let
-then contradicts in the lower part, and in the upper part, contradicts and in the upper part, contradicts in the lower part.
-Obviously, this may be generalized to paths.
-Consider Yablo’s original construction:
-Example 1.5
-Let the nodes be and the arrows for expressed as a relation by and as a logical formula by
-Thus For any we have a contradiction by and for any and for any for a suitable
-Example 1.6
-This Example shows that infinitely many finitely branching points cannot always replace infinite branching - there is an infinite “procrastination branch” or “escape branch”.
-This modification of the Yablo structure has one acceptable valuation for
-Let as usual, and introduce new
-Let with
-If then by so, generally,
-if then and
-If then so if etc., so, generally,
-if then or
-Suppose now then for all and for all By there is and a contradiction, or for all again a contradiction.
-But is possible, by setting and for all
-Thus, replacing infinite branching by an infinite number of finite branching does not work for the Yablo construction, as we can always chose the “procrastinating” branch.
-Diagram 1.1
-Escape Path
-Remark 1.5
-For a representation result, we need enough paths in the graph considered to form contradictions.
-This is not trivial.
-Suppose that and are both negative, and is the only path from to then the condition is obviously false.
-The author does not know how to characterize graphs which satisfy the condition directly.
-Some kind of “richness” for the graph will probably have to hold.
-(This is covered in Yablo’s construction by transitivity.)
-1.2.4 Interplay of the Graph and the Logical Side
-We can either think on the logical level with formulas, or graphically with conflicting paths, as in the following Fact.
-We need infinite depth and width in our constructions:
-Fact 1.6
-The construction needs infinite depth,
-the logic as used in Yablo’s construction is not compact,
-the construction needs infinite width, i.e. the logic cannot be classical.
-Proof
-Let be a minimal element, then we can chose an arbitrary truth value and propagate this value upwards. If there are no infinite descending chains, we can do this for the whole construction.
-The logic as used in Yablo’s construction is not compact:
-Trivial. (Take This is obviously inconsistent, but no finite subset is.).
-It is impossible to construct a Yablo-like structure with classical logic:
-Take an acyclic graph, and interpret it as in Yablo’s construction.
-Wlog., we may assume the graph is connected.
-Suppose it shows that cannot be given a truth value.
-Then the set of formulas showing this does not have a model, so it is inconsistent.
-If the formulas were classical, it would have a finite, inconsistent subset, Define the depth of a formula as the shortest path from to this formula.
-There is a (finite) such that all formulas in have depth Give all formulas of depth (arbitrary) truth values, and work upwards using truth functions.
-As the graph is acyclic, this is possible.
-Finally, has a truth value.
-Thus, we need the infinite
-The following example illustrates (partially) unfinished contradictions, and resulting escape possibilities.
-Example 1.7
-There is a fundamental difference between the contradictions (for and
-Consider the case The path is blocked at So we cannot go to by two contradicting possibilities.
-The path is, however, not blocked at (or elsewhere), so we have two contradicting possibilities to go from to This basically behaves like and we have an escape possibility.
-So, if we try to append the same construction at we have an escape possibility, always making the origin of the construction negative.
-This is not true for We can append the same construction at as will always be at the opposite polarity of due to , the alternative path ends at so there is no escape.
-So a negative start at the original diagram leads to a positive start in the appended diagram.
-Diagram 1.2
-Diagram Postponed Contradictions
-1.2.5 A Third Truth Value
-See also Section 7.1.2, Fact 7.1.1 in [Sch22] for a less systematic approach.
-Remark 1.7
-We consider a third truth value, in addition to and
-means for a formula that neither nor are true (in the given set of formulas). Of course, this is impossible in a classical framework.
-Let be the truth value of
-Let be an abbreviation for and
-Let
-We have
-analogously.
-Likewise
-We thus have:
-and
-Using the operations for truth values
-and above results, we have
-(or and thus the order
-(and finally the inverse operations
 1.3 Strategy
 These comments will be clearer in hindsight, after having read the details. Still the best place to put them seems to be in general remarks, and go into details later on.
-The comments on “ is false ” show that there is no other way to express this in an acyclic graph of our type.
-The comments on the construction in the limit solve a seeming contradiction by chosing the right level of abstraction, i.e. by arguing on the level of instances.
-1.3.1 Expressing “x is false” In an Acyclic Graph
-In an acyclic graph, we cannot express something like we have to use different means to express that is impossible.
-As we can always fill in truth values (including from the bottom, we cannot force by looking upwards of (or sideways, i.e. there is no monotone path connecting with some So we have to look downward.
-We may use or if we have the constant We will see later, that this not interesting, see Definition 2.1 (page 2.1) and Remark 2.1 (page 2.1).
-This is a contradiction made of (graphically speaking) two paths of length one each.
-Interesting contradictions are made of slightly more complicated paths, like the Yablo Cell meaning
-Basically, the only way to express that is impossible, is by the logical formula
-A slightly different argument: How can we express that is wrong in a finite acyclic graph?
-Consider the set of minimal elements below , and the formula corresponding to the graph in DNF of these elements.
-This might be
-(the graph is trivial, just
-or a tautology or contradiction.
-a disjunction of conjunctions of elements in , or their negations. By acyclicity,
-This argument is not totally satisfactory. For this reason, we reformulate:
-We examine graphs, where there existe a node such that as well as -x will lead to a contradiction.
+Definition 1.5
+Conditions (C1) and (C2)
+We sometimes write when we want to emphasize that we assume is positive, similarly for the negative case.
+Consider in the Yablo construction (but this applies to all
+We denote by (C1) (for the condition that has to be contradictory, and by (C2) the condition that has to be contradictory, which means by that all have to be contradictory, roughly, in graph language, that all paths from have to lead to a contradiction.
 1.3.1 Satisfying the Antagonistic Requirements (C1) and (C2) In the Limit
@@
 We proceed inductively as Yablo did, but we also see that we basically have no other choice (modulo minor modifications).
-So we want to express by an infinite graph that both and are inconsistent.
-As we saw in Section 1.3.1 (page 1.3.1), this must be expressed by suitable contradictions.
-Therefore, we treat such contradictions as basic building blocks, as Yablo did.
+So we want to express by an infinite graph that both and are inconsistent. See Definition 1.2 (page 1.2). Therefore, we treat such contradictions as basic building blocks, as Yablo did.
 The problems are then
@@
 So we alternate between and a tautology, and it is not clear what such constructions are in the limit, when we leave classical logic.
-This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 3 (page 3).
-A summary: we have to look at instances.
+This view is too abstract, we have to look at details, as they are given in the proof of validity for Yablo’s construction, or our reconstruction in Section 3 (page 3), we have to look at instances.
 - (2.1)
@@
 So we might first treat these two branches ((C1) problems) before going back to the new (C2) problem created by the first new branch.)
-1.4 The “Right” Level of Abstraction
-1.4.1 Introduction
-The basic elements in Yablo’s construction are negative arrows, from which Yablo Cells are built.
-We show here that we can build arbitrarily complex structures equivalent to negative arrows, and as a matter of fact to any propositional logical operator.
-This suggests that the right level of abstraction to consider more general Yablo-like structures is not the level of single arrows, but rather of paths.
-Remark 1.8
-In the diagrams in Diagram 1.3 (page 1.3) the inner path is barred by , and the outer path contains 3 or 2 negations.
+Example 1.1
+This Example shows that infinitely many finitely branching points cannot always replace infinite branching - there is an infinite “escape branch” or path.
+See below, Definition 1.6 (page 1.6).
+This modification of the Yablo structure has one acceptable valuation for
+Let as usual, and introduce new
+Let with
+If then by so, generally,
+if then and
+If then so if etc., so, generally,
+if then or
+Suppose now then for all and for all By there is and a contradiction, or for all again a contradiction.
+But is possible, by setting and for all
+Thus, replacing infinite branching by an infinite number of finite branching does not work for the Yablo construction, as we can always chose the “escape” branch, see below.
 Definition 1.6
-We define negations and their type.
-A negation (diagram) is a diagram all of whose arrows are of the kind.
-We define negation diagrams, or, simply, negations, and their type.
-An arrow is a negation of type 0.
-all negations inside are of type and at least one negation inside is of type
-Remark 1.9
-Thus, it seems very difficult to describe Yablo type diagrams on the level of single arrows. It seems to be the wrong level of abstraction.
-We now give some examples.
-1.4.2 Some Examples
-Example 1.8
-Negation (1)
-so and
-Negation (2)
-so and continue as above.
-Identity
-so and
-TRUE
-Thus, and
-so and
-Diagram 1.3
-Propositional Formulas 1
-Diagram 1.4
-Propositional Formulas 2
-1.5 Prerequisites
-We work with pure conjunctions of positive or negative propositional variables, like except for the following:
-we may use a constant TRUE or
-we may use formulas of the type
-so the following are also permitted: etc.
-This is used only to “neutralize” unwanted parts of the graph we characterize, see Fact 4.4 (page 4.4).
-- (2)
-- (3)
-- (4)
+When we try to make a structure contradictory for some value of the origin, but find a way (a valuation) to avoid or escape a contradiction, we call such a valuation an escape path, branch, or possibilty.
+This is then a model consistent with the theory of the graph.
+This seems an intuitive description how the structure fails to be contradictory for this valuation.
 2 Elementary Cells
 2.1 Introduction
-Elementary cells have to be simple constructions providing contradictions (in some cases, e.g.
-when the origin, usually called or so, cannot have a certain truth value, as this will lead, within the cell, to a contradiction).
-Thus
-They must be contradictory, and thus exclude at least one instance.
-In addition, they should be simple, e.g. they will not contain a sequence or so without branching, as we could replace this with
-(We will see later how to relinquish this requirement.)
-More importantly, we must be able to construct a Yablo-like structure, built up by this type of elementary cell only (and perhaps some simpler auxiliary structures).
-They must serve as building blocks for a suitable stucture.
-This prevents “cheating”, letting other constructions do the crucial work.
-- (4)
+We try to generalize Yablo cells to some other types of cells.
+(C1) They will have the form of two paths which branch say at and meet again at (but not before), and are contradictory.
+(C2) Both branches will have to lead to new contradiction cells if
+We require that it is possible to construct a structure with such cells starting, say at such that and are contradictory.
+In addition, we require that this is possible using only the contradiction cells of this type. This prevents “cheating”, where the real work is done by some other type of cell.
+We will impose syntactical limitations on the formulas attached to the nodes in the cells.
 2.2 Basic Discussion
-See Section 7.3 in [Sch22] for more material.
-We try to describe here the basic constructions of contradictions.
-As said in Fact 1.1 (page 1.1) (and re-written graph-theoretically), these consist of two branches with common origin, which meet again, and have different polarity. We call such constructions cells.
-First, we want to exclude some trivialities. We describe only the part beginning at the branching point, not before, not after.
-Second, we will define a simple hierarchy of such cells, and will allow a cell only the use of simpler cells as substructure.
-This prevents “cheating”.
-(See Section 2.1 (page 2.1).) Thus, we look at those cells only which allow to construct a Yablo-like structure without the use of more complicate cells.
-We first use (almost) only negative arrows, and nodes whose formulas are conjunctions. We will see how to generalize to more complicated paths.
 Definition 2.1
@@
 Remark 2.1
-An escape problem with for the trivial contradiction, so (See Definition 2.1 (page 2.1).) If we continue with so etc. we have an escape sequence etc., which never meets a contradiction.
-Thus, is not a contradiction cell for our purposes, has a model defined by the escape sequence.
+- (1)
 Suppose we have an escape problem, like thus
@@
 This, however, is not possible.
-There are, whatever the inner structure, just four possibilities: it might be equivalent to The first two cases are trivial. Suppose Then Suppose Then
+There are, whatever the inner structure, just four possibilities: it might be equivalent to
+The first two cases are trivial, as we have the possibilities or so is always a possibility.
+Suppose Then
+Suppose Then
 In general, suppose we have a partial graph with origin and end point (among perhaps other end points), and, say results in choices or (due to internal choices by OR).
-Example 2.1
-We consider now some simple, contradictory cells.
-They should not only be contradictory for the case but also be a potential start for the case without using more complex cells.
-(Otherwise, we postpone the solution, and may forget the overly simple start.)
-We will consider in Section 2.5 (page 2.5) a more complicated contradiction cell (“Diamond”), and discuss here only simple cells.
+If we continue with a copy of now grafting on etc., we may not achieve a desired result (whereas a different may achieve it).
+Remark 2.2
+We consider here cells with three arrows.
 Note that the following considerations apply also to cells formed by paths consisting of more than one arrow, and not only to cells formed by single arrows.
@@
 In particular, also in more complicated cells all sides have to negative.
-The cell with 2 arrows.
-It corresponds to the formula graphically, it has a positive and a negative arrow from to so exactly one of and is negative.
-If is positive, we have a contradiction.
-Of course, appending at a Yablo Cell (see below, case (2.3)) may be the beginning of a contradictory structure, but this is “cheating”, we use a more complex cell.
-Cells with 3 arrows.
-In the graph with not annotated arrows, they have the form -z, -z, where the lines may stand for or
+In the graph with not annotated arrows, they have the form , , where the lines may stand for or
 Again, we want a contradiction for positive, so we need an at For the same reason, the number of negative arrows may not be even. So assume the number of negative arrows is one or three.
@@
 This will be discussed in detail in the rest of the paper.
-But we see already that both paths, and change sign, so and will be positive, appending the same type of cell at and solves the problem (locally), and offers no escape.
-We call sometimes the head of the Yablo cell, its foot, and its knee.
-If we combine Yablo Cells, the knee for one cell may become the head for another Cell, etc.
+But we see already that both paths, and change sign, so and will be positive when is negative, appending the same type of cell at and solves the problem (locally), and offers no escape.
 One negative arrow:
@@
 Consider the case
-Upper part:
-Lower part:
+Consider the theory indicates an escape path.
+Consider the theory indicates an escape path.
 Consider the case
+Consider the theory indicates an escape path.
 Consider the case
-Here,
-Diagram 2.1
-Simple Contradictions
-Diagram 2.2
-Diagram ESC-1
-Diagram 2.3
-Diagram ESC-2
-Diagram 2.4
-Diagram ESC-3
+Consider the theory indicates an escape path.
+See also the discussion of “Saw Blades” in [Sch22], Section 3.
+2.3 Values for y and z in Yablo Cells
+We consider here what happens when we give and truth values In addition, we consider the variants and
 Remark 2.3
-The distinction between and i.e.
-between knee and foot, is conceptually important.
-In the case we have at a complete contradiction, at we have not yet constructed a contradiction.
-Thus, if we have at again an (as at this becomes an and we have to construct a contradiction for all (or not only for Otherwise, we have an escape possibility for Obviously, the contradiction need not be immediate at the important property is that ALL paths through all lead to a contradiction, and the simplest way is to have the contradiction immediately at - as in Yablo’s construction, and our saw blades.
-See also the discussion of “Saw Blades” in [Sch22], Section 3.
-As we work for a contradiction in the case x-, too, the simplest way to achieve this is to have a negative arrow and at again an This gives a chance to construct a contradiction at Of course, we have to construct a contradiction at too, as in the case we have an at
-Of course, we may have branches originating at which all lead to contradictions in the case so (resp. at is possible, too.
-But, for the simple construction, we need between and and at And this leads to the construction of contradictions for all (or as just mentioned above.
-2.3 Values for y and z in Yablo Cells
-We consider here what happens when we give and values In addition, we consider the variants and
-Remark 2.3
-False, True
 We present a systematic treatment of variants of the “environment” of the Yablo triangle (and the diamond).
@@
 Do we obtain
-- (2.4)
+Do we obtain infinite width and depth by suitable combinations of cells of the same type, see Section 2.1 (page 2.1), and [Sch25], Fact 2.4 or [Sch23b], Fact 1.4.
+See also Section 3 (page 3).
+This requires continuation at both and with constructions of value case See below (4.3) for an (unimportant) exception.
 We consider the cases in all possible combinations:
@@
 Thus, we look at e.g. case i.e. and etc., and consider whether the requirements in (2) above are satisfied.
-We first look only at the minimal requirements
-has to be contradictory, requirement (2.1) above.
-must not be classical (i.e. not or requirement (2.3) above.
-The cases:
-(2) (requirement (2.3))
-Consider so for any so these cases are classical.
-Consider For is classical, so is, too.
-so
-Thus, is the only non-classical case among
-Consider In cases so too.
-In case so
-In case so
-So is the only classical case among the
-So the interesting cases left are
-(1) (requirement (2.1))
-We now check for local contradiction in the case i.e. must not be
-may only be iff But if then only in case
-Next, we have to look at and see if every path from leads to a contradiction, requirement (2.2) above. So neither nor must be
-So and will not work.
-Consider i.e.
-Then in and so this does not work.
-so this works.
-so this works.
-Consider i.e.
-Then or and all cases work, except
-In summary: in exactly the following cases: so they are excluded.
-We summarize the results for the three requirements above:
+We look at above requirements in the following order: (2.3), (2.1), (2.2). This is a tedious exercise and left to the reader, these are the results:
 (2.3): are ok.
@@
 In summary, only satisfy all three requirements, (2.1), (2.2), (2.3).
-We look at these cases.
+We look at the results in the interesting cases.
 Here, and in (3.6), stands for which is of type likewise as we will append in the full structure at and a construction with value
@@
 Thus,
-Thus,
-identical to
-Thus,
+so
+(Yablo) Thus,
+identical to Thus,
 Thus,
@@
 Summary of the (in this context) important constructions
+- (4.1)
+- (4.2)
+The simplified Saw Blade tooth is of the type with type above, see Section 7.6 in [Sch22], or Section 3 in [Sch23b].
+(In this construction, will be (except for the start) again of type as we obtain this by transitivity, forced by condition (C1), see Construction 3.1 (page 3.1), and Diagram 3.1 (page 3.1), steps (2.3), (2.5), (2.6)).
+Both the original Saw Blade and the simplified one contain the Yablo construction.
+See, e.g., Diagram 3.1, upper part, in [Sch23b]: has the role of in Yablo’s construction.
+is a Yablo cell.
+must be the head of a new Yablo cell, starting e.g.
+with has to be contradicted, so we need so must be the head of a new Yablo cell, so we have some etc.
+- (4.4)
 2.4 Yablo Cells Formed By Paths
@@
 2.4.1 Branching points
-We have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).
+Remark 2.4
+is the origin of a new cell, so we have to branch at (Anyway, we have to branch on the path , otherwise, we have the trivial contradiction (and an escape possibility).)
 Branching at
@@
 We may branch on the already existing oder , and continue to build a (finite) contradiction to
-This however, is possible only a finite number of times, e.g. then etc., see Example 1.6 (page 1.6). This constructs an infinite sequence of choices which offers an escape possibility.
+This however, is possible only a finite number of times, e.g. then etc., see Example 1.1 (page 1.1). This constructs an infinite sequence of choices which offers an escape possibility.
 Thus we need infinite branching (and non-classical logic).
@@
 By full transitivity of the final construction, any infinite subsequence of the construction has the same properties as the full construction.
+Branching at on the path :
+Consider the case Whatever the branching situation on is (just one branching, or multiple branchings), for the branch considered in , the path is supposed to create a contradiction.
+Suppose we have a branching at where i.e.
+it is an “OR”, then we may decide at otherwise, and not have a contradiction with the branch .
+Thus, we first chose a path via and then chose a contradicting path (or directly ).
+In Yablo’s construction, this is trivial, as is a direct link.
+Definition 2.2
+If , then there is forming a Yablo cell, i.e. all three components , , are negative, such that contains no branching point with for
+We call this weak transitivity.
 2.4.2 Paths
@@
 We construct for as for , i.e.
+Note: need not be contradictory for , as is contradictory, thus, will be contradictory ( will be AND for ), so (C2) is satisfied.
+Thus, has the role of and we may invoke our rule that contradictions should happen immediately, and not be postponed.
 Case 2: AND
@@
 We need, however, that all branches lead to contradictions (condition (C2)).
+We might identify and , and impose this as an additional requirement for the structures we examine, along with “no postponement of contradictions”.
 We branch on , and have , .
@@
 Consider e.g. , , .
-As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (” pipeline “), we may assume for simplicity that we always start contradictions at
-Diagram 2.5
+As branching at is not different from branching before or after we assume for simplicity that we branch at Moreover, as argued above (“pipeline”), we may assume for simplicity that we always start contradictions at
 2.5 The Problem with Diamonds
@@
 We have a conflict between the diamonds starting at and and the diamond starting at
-If then y- and As the choices at and are independent, any branch -z, combined with any branch -z, must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
-However, if is positive, the diamond -z, has to be contradictory, so not both branches may be negative.
-A solution is to “synchronise” the choices at and which can be done e.g. by the formula
-and This is a different type of formula, moreoever corresponding arrows from are missing.
-Even if we do not consider the paths for the diamond starting at we have and so we can make and false, and chose and which results in an escape possibility (we have to consider new diamonds starting at etc.)
-This formula has an additional flaw, we want to speak about paths we chose, and not just their end points.
-This can be done with introducing additional points on the arrows, e.g.
-the arrow is replaced by now we can speak about the arrow we have given it a name.
-We may also put such information in a background theory, which need no be negated.
-But all this leads too far from the basically simple formalism of Yablo’s construction, so we will not discuss this any further.
+If then and .
+As the choices at and are independent, any branch , , combined with any branch , , must be conflicting, thus, given is negative, all branches on the left must be negative, likewise, all branches on the right must positive.
+However, if is positive, the diamond , has to be contradictory, so not both branches may be negative, likewise for
+A solution is to “synchronise” the choices at and Informally, this is possible by: “ and at choose the arrows or and at choose the arrows or ”.
+However we express this, the negation is “ or choose at or choose at the arrow ”, which has nothing to do with the usual negations in Yablo structures.
+A way out might be to put the choices at and into a background theory which is not affected by negation.
+Even if we consider arbitrary DNF formulas, we cannot speak about choices at and or arrows like Thus, within our framework, the diamond is NOT a suitable contradiction cell.
 (Yablo’s construction does not need synchronisation, this is done automatically by the universal quantifier.)
 The simplified version with recursion on one side only, here on the left only, is nothing but the Yablo triangle:
-Thus, we conclude that the original version with 4 points and full recursion is beyond our scope, and the simplified version (see above) is nothing but the Yablo triangle.
-Diagram 2.6
+Diagram 2.1
 Nested Diamonds, Details
-Diagram 2.7
-Diamond Essentials, Recursion on the Left, Simple Contradiction on the Right
 3 Combining Cells
-3.1 Overview
-We follow Yablo’s construction.
-(C1): We have to make contradictory.
-We first try to make “directly” contradictory. That is, we will not first go to and make contradictory, nor will we make a disjunction at and then make each disjunct contradictory.
-(This is not an important simplification, as more complicated constructions will finally use our simple one, too. See e.g. Diagram 1.2 (page 1.2) and its discussion.)
-As Yablo did, we set and
-The first step is done, is contradictory. Probabilistically, of the possibilities via are contradictory.
-(C2) We have to make contradictory.
-Now, must be contradictory, too. As we do not know which one is positive, we must make both and contradictory - just as a disjunction is false if every disjunct is false.
-Consider, e.g., Again, we will make directly contradictory (as above for so is a conjunction, say (we might re-use here, but this will not be important) and add e.g.
-and to for the contradiction, so we have
-But now, we have destroyed the contradiction at as for is a disjunction, and have to make ALL possibilities for contradictory for so far only the possibility is contradictory.
-Probabilistically, of the possibilities at are open.
-To create a contradiction for the possibility we introduce new contradictions by adding to so
-Etc.
-See below and Section 3.3.3. in [Sch23b].
-Note that for (C1), i.e. for and for (C2), i.e. for we have to make for any disjunction all possibilities contradictory, and for any conjunction it suffices to make one conjunct contradictory, as and
-Of course, following a path in the graph, conjunctions for will become disjunctions for one of the confusing aspects of reasoning.
-(C1) and (C2) are antagonistic requirements.
-Satisfying (C1) creates new problems for (C2), and vice versa.
-In the limit, all requirements are satisfied, since all problems will solved in the next step, at the price of creating new problems, which will be satisfied in the next step.
-Note that we don’t approximate, as long we are finite, we always have only in the infinite case we have
-3.2 In More Detail
-We now consider the inductive construction of the Yablo structure, see Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2)
-In (2) and (5), we have unrelated points and we may consider the model sets to be orthogonal. - We only indicate model sets briefly, without going into details (U will stand for the universe).
-abbreviated
-has to lead to a contradiction for
-preparation,
-was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
-finish,
-(branch has to be contradicted for add
-has to lead to a contradiction for
-preparation,
-finish,
-(branch has to be contradicted for add
-(branch has to be contradicted for add
+We follow Yablo’s construction, and refer the reader to Section 1.3.1 (page 1.3.1) for an abstract discussion of this construction.
 Construction 3.1
-Summary:
+The principle
+If has to be the start of a Yablo cell, but is a terminal node, add an arrow
+If has to be the start of a Yablo cell, exists, but not add (transitivity).
+If will be the start of a new Yablo cell, if
+Details
+thus
+will be the start of new Yablo cells.
+By (C2), has to lead to a contradiction for
+preparation: we use as part of the new Yablo cell, and add the new arrow as second part of the new cell, where is a new point.
+finish: We add the new arrow as third part of the new cell.
+Note: was (part of) a full contradiction, this is now a partial contradiction, as the new possibility is added.
+By (C1),
+(branch has to be contradicted for add thus
+will be the start of a new Yablo cell.
+By (C2), has to lead to a contradiction for
+preparation: we use as part of the new Yablo cell, and add the new arrow as second part of the new cell, where is a new point.
+finish: We add the new arrow as third part of the new cell.
+By (C1),
+(branch has to be contradicted for add Thus
+Complete Yablo cell starting at
+By (C1),
+(branch has to be contradicted for add Thus
+will be the start of a new Yablo cell.
+Note:
+Summary (in the limit):
+Main properties:
 Thus, there are arrows for all and the construction is transitive for the
@@
 Thus, every arrow from any to any goes to the head of a Yablo cell, and not only the arrows from
-This property is “accidental”, and due to the fact that for any arrow there is also an arrow and property (2) holds for by prerequisite.
+This property is “accidental”, and due to the fact that for any arrow there is also an arrow and (C2) holds for by prerequisite.
 The construction has infinite depth and branching.
@@
 Our construction is somewhat arbitrary. We construct a contradiction for by and use the arrow as part of the contradiction for This, of course, is a special case.
-Consider Diagram 3.3 (page 3.3). Here, the contradiction for is as above, but we begin a totally new contradiction for likewise for etc.
-But, we see, the old construction of Diagram 3.1 (page 3.1) and Diagram 3.2 (page 3.2) is part of the more general construction of Diagram 3.3 (page 3.3), so the old construction is really minimal.
+We may begin a totally new contradiction for likewise for etc.
+But the old construction of Diagram 3.1 (page 3.1) is part of this somewhat more general construction, so the old construction is really minimal.
 We have to be a bit careful here. If we set etc. then we have etc. Instead, we may set so we have as desired, or we append at recursively the same construction - as in “Saw Blades”.
+Remark 3.2
+Above construction can be read differently: Suppose we have an (annotated) graph such that a node has truth value Then we can do above construction.
+E.g., has to be contradictory, (C1), so have two contradictory paths from to, say and some on the path Both paths have to lead to contradictions for (C2).
+Exactly one of them has to be on one of the paths (and not at the end), say this is (if both are on the paths, we would have the diamond).
+This way, we can reconstruct the Yablo construction within Of course, there may be additional branchings, as discussed in Section 2.4 (page 2.4).
+Next, we show how to neutralise them.
+Definition 3.1
+Consider an annotated graph a subset of its nodes, and the subgraph generated by in Consider now where in is identical to in restricted to the nodes in
+Remark 3.3
+above is identical up to logical equivalence for the nodes in to the following modification of replace for all and arbitrary, if and is terminal by and if is not terminal, by We call tis the isolation of in
+Example: Let be as follows: and then
+Consequently, in Remark 3.2 (page 3.2), the isolation is equivalent to the original Yablo structure.
 Diagram 3.1
-Diagram Inductive Construction 1
-Diagram 3.2
-Diagram Inductive Construction 2
-Diagram 3.3
-Diagram Inductive Construction Core
+Diagram Inductive Construction
 4 Result
-Consider structures of the type (like ’s), with one small addition: we also allow formulas of the type either as (stand alone), or as one of several elements of a conjunction where is of the type It is important that these additions also cooperate with our truth value
-(This is more general than adding the constant TRUE, as it permits chaining of the new formulas, but we also add TRUE for terminal nodes.)
-In addition, we assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it.
+4.1 Prerequisites (Limitations)
+We assume that problems are attacked immediately: If we need a contradiction in the construction, we add it, and not a path leading to it.
 We also assume we use only one type of contradiction cell, and work not first with unsuitable cells, postponing the use of efficient cells.
@@
 See here e.g.
-the beginning of Section 2.2 (page 2.2), and Remark 2.1 (page 2.1), and Example 1.7 (page 1.7), with Diagram 1.2 (page 1.2).
+the beginning of Section 2.2 (page 2.2), and Remark 2.1 (page 2.1).
 Without these assumptions, we have to consider many complications, which, however, do not present deeper insights.
-4.1 A Suitable Injection Does the Job of Characterisation
+Given these limitations, we show that Yablo’s structure is minimal for acyclic structures showing that and lead to contradictions, and that any unlabelled, directed graph permits a labelling showing the same iff Yablo’s original structure has a suitable injection into the graph in question.
+4.2 A Suitable Injection From the Yablo Structure
 Consider Yablo’s original structure YS, and some directed acyclic graph YS’.
-Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory iff the following conditions hold.
+Then there is an interpretation of YS’ in the extended language as described above which makes some element (which we might make initial) contradictory, as well as contradictory if the following conditions hold.
 Condition 4.1
-There is an injection from YS to YS’, more precisely from the nodes of YS to the nodes of YS’, s.t., if there is a path from to in YS, then there is a path from to
-For THIS injection, there is a valuation of the arrows in (the other arrows are not interesting) s.t., if form a contradiction in YS, then likewise for (Contradictions are preserved under
-4.2 Proof
-Fact 4.2
-Yablo’s structure is the simplest structure showing that both and are contradictory.
+There is an injection from YS to YS’, more precisely from the nodes of YS to the nodes of YS’, s.t., if there is a path from to in YS, then there is a path from to in YS’, in particular, if is a direct successor of on such a path, there must be an arrow in YS’.
+For this injection, we need the following condition
+Let be the subgraph consisting of all nodes and the intermediate nodes on these paths.
+We have to be able to construct a partial valuation in YS’, defined on the nodes in YS” (of course respecting the arrows in s.t.
+for any two nodes if there is an arrow in YS, then iff This need not hold for intermediate nodes.
+Thus, contradictions are preserved for the paths in YS’.
+Definition 4.1
+We extend this partial valuation in YS’ to a total valuation in YS’ as follows:
+If is a node in set there is an arrow in Thus, we respect the arrows in YS’, and assures that the result is and not if
+If is a node in YS”: We have already a formula describing the behaviour for all successors of in YS”. We now have to respect the other arrows in YS’, and set
+there is an arrow in YS’, and not a successor of in above paths
+Remark 4.1
+- (1)
+The second condition expresses “richness”, that we have enough independent arrows.
+4.3 For the converse
+Fact 4.3
+Yablo’s structure is the simplest structure showing that both and are contradictory, and contained in all such structures.
 Proof
@@
 - (3)
-Remark 4.3
-Consider
-We may chain such formulas together, like etc., which is impossible for the constant TRUE.
-is preserved, if then too. Likewise, if and then
-Consider Yablo adding
-EG: so , this works well classically, but also for as
-We may also continue beyond : or then too.
-If we have mixed chains of and the last element will decide the result (trivial).
 Fact 4.4
-Consider (without the other elements etc. in
-Now consider the full YS’, and attribute values to the nodes as follows, respecting the values already given in (1):
-If is minimal, then
-Otherwise:
-If dann
-arrow
-Note that these arrows have no value, i.e. we have no distinction between and
-If then
-: arrow
-: arrow
-: arrow
-Note that arrows inside have a value, these values have to be preserved. The other arrows have no value, and we treat them as positive.
-Moreover, we did not distinguish between minimal and not minimal this does not matter. If we want to distinguish we have to differentiate the last conjunct further.
-- (3)
-Remark 4.5
-Asssume for simplicity that (arrows may be negative, too), likewise for etc.
-We have paths etc., assume for simplicity that the only new intermediate node is respectively).
-Let is not in
-We set arbitrary, Thus, so and so
-Of course, instead of going through the additional path might also go through etc. This changes nothing. We have neutralised the path not in
-Diagram 4.1
-Diagram
-Diagram 4.2
-Diagram
+Suppose the graph has a valuation with an element
+Proof
+By Fact 4.3 (page 4.3), or directly by the construction Construction 3.1 (page 3.1) INSIDE we have an analogue of the Yablo structure in perhaps with paths instead of simple arrows, and perhaps with many unnecessary parts.
+We can eliminate the unnecessary parts as follows, (for ease of reading with a simple example, analogously to Definition 4.1 (page 4.1)):
+Suppose we have where the part is unnecessary. Then replace by so we have
+Suppose we have where is unnecessary, then replace by so we have
+As we have DNF formulas, this takes care of it.
+5 Acknowledgements
+The author would like to thank the anonymous referee for his many valuable suggestions, and the editorial board and Ms. Jane Spurr for their help.